### v1

# Blind Inverse Problem Solving Made Easy by Text-to-Image Latent Diffusion

Abstract
Blind inverse problems, where both the target data and forward operator are unknown, are crucial to many computer vision applications. Existing methods often depend on restrictive assumptions such as additional training, operator linearity, or narrow image distributions, thus limiting their generalizability. In this work, we present LADiBI, a training-free framework that uses large-scale text-to-image diffusion models to solve blind inverse problems with minimal assumptions. By leveraging natural language prompts, LADiBI jointly models priors for both the target image and operator, allowing for flexible adaptation across a variety of tasks. Additionally, we propose a novel posterior sampling approach that combines effective operator initialization with iterative refinement, enabling LADiBI to operate without predefined operator forms. Our experiments show that LADiBI is capable of solving a broad range of image restoration tasks, including both linear and nonlinear problems, on diverse target image distributions.
Figure 1: Our proposed LADiBI is a training-free blind inverse problem solving algorithm using large pre-trained text-to-image diffusion models. LADiBI is applicable to a wide variety of image distribution as well as operators with minimal modeling assumptions imposed.
1 Introduction
Inverse problems are fundamental to many computer vision tasks, where the goal is to recover unknown data ${\bm{x}}$ from observed measurements ${\bm{y}}$. Formally, these problems can be expressed as ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$, where ${\mathcal{A}}$ is a operator representing the forward process parametrized by $\phi$, and ${\bm{n}}$ is the measurement noise. These problems are critical in fields such as medical imaging, computational photography and signal processing, as they are inherently tied to real-world scenarios like image decompression, deblurring, and super-resolution  [41, 43, 13, 17]. Recently, deep learning-based methods  [30, 22, 44, 8, 7, 10] have emerged as powerful tools for solving inverse problems. In particular, diffusion models offer an elegant solution that incorporates pre-trained models in diverse conditional settings, providing flexible and scalable frameworks for reconstructing high-quality data by guiding the generation process through the measurement constraints [19, 36, 3, 37, 42, 14].
However, most of these efforts focus on inverse problems where the operator ${\mathcal{A}}_{\phi}$ is assumed to be known. In practice, the forward operator is often unknown, leading to what are termed blind inverse problems. Solving these blind inverse problems remains a significant challenge due to their ill-posed nature. Consequently, current methods address this problem by introducing various assumptions about the prior distributions of the target data and unknown operator. These assumptions are typically incorporated through a combination of the following strategies: (1) using explicit formulas or specific neural network architectures to hand-craft inductive biases, (2) simplifying the operator with linear assumptions, (3) restricting target distributions to narrow image classes, (4) collecting measurement, operator or image datasets to train models tailored to specific tasks. While effective in certain cases, these solutions often entail significant computational and manual overhead, limited flexibility, and substantial barriers to practical deployment due to complex model training, costly data curation, unrealistic assumptions, and laborious hyperparameter tuning. These limitations raise a critical question: can we design an algorithm that applies to diverse operators and target image distributions without additional training or data collection?
To tackle this challenge, we formulate this problem as a Bayesian inference problem, where the optimal solution is to sample from the posterior $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi})$. State-of-the-art diffusion-based methods often simplify this problem by assuming independence between the target image and operator distributions [8, 27, 23, 34]. Although this independence assumption reduces the modeling difficulty, we challenge its validity by observing the correlation between the target image distribution and the operator distribution, and argue that this assumption heavily restricted the flexibility of the current diffusion-based approaches. To address this problem effectively, we first propose to estimate the joint distribution $p({\bm{x}},{\mathcal{A}}_{\phi})$ instead. While this estimation seems intractable on the first glance, there actually exists a surprisingly simple solution – large-scale pre-trained text-to-image diffusion models already capture diverse target data and measurement distributions. Moreover, many inverse problems can be described using natural language, allowing tasks like deblurring or super-resolution to be expressed through prompts such as “high-definition, clear image” for the target and “blurry, low-quality” for the measurement. By simply prompting with classifier-free guidance [15], we can approximate the score of this challenging joint prior distributions across a wide range of image and operator types using the same base model, significantly extending the flexibility of the current diffusion-based framework.
In addition to a strong prior, an effective posterior sampling technique is also crucial for meeting measurement constraints. Existing methods generally fall into two categories: Gibbs or EM-style alternating optimization [27, 23], which suits diffusion-based setups without additional training but is sensitive to operator initialization and tuning, and a two-step approach that first estimates the operator followed by solving the non-blind problem [34], which often requires training and strong assumptions. To obtain the best of both worlds, we propose a hybrid approach: first, we obtain a reasonable initial estimate of the operator parameters, and then apply the alternating optimization to iteratively refine both operator parameters and data estimates throughout the diffusion process. Obtaining a reliable initial operator estimate can be difficult for broad functional classes like neural networks. To address this, we introduce a novel initialization scheme that leverages the pseudo-supervision signals from multiple lower-quality data estimation generated by fast posterior diffusion sampling. This approach eliminates the assumption of a specific form of operators, allowing for nonlinear blind inverse problem solving and flexible operator parametrization.
Combining the strong prior with the effective posterior sampling technique, we introduce Language-Assisted Diffusion for Blind Inverse problems (LADiBI), a training-free method that leverages large-scale text-to-image diffusion models to solve a broad range of blind inverse problems with minimal assumptions. LADiBI can be directly applied across diverse data distributions and allows for easy specification of task-specific modeling assumptions through simple prompt adjustments. Algorithm 1 provides an overview of LADiBI, which can be easily adapted from the standard inference algorithm used in popular text-to-image diffusion models. Unlike existing methods, LADiBI requires no model retraining or reselection for different target data distributions or operator functions; instead, all prior parameterization is encoded directly in the prompt, which users can adjust as needed. Notably, we do not assume linearity of the operator, making LADiBI, to the best of our knowledge, the most generalizable approach to blind inverse problem solving in image restoration.
We evaluate LADiBI against state-of-the-art baselines across several blind image restoration tasks, including two linear inverse problems (motion deblurring and Gaussian deblurring) and a nonlinear inverse problem (JPEG decompression), on various image distributions, as illustrated in Figure 1. In the linear cases, our method matches the performance of the state-of-the-art approaches, which require extensive assumptions and model specifications. Additionally, LADiBI is the only method tested that successfully performs JPEG decompression without any information about the task, such as the compression algorithm, the quantization table or factors, relying solely on observations of the compressed images.
2 Background & Related Works
Table 1: A conceptual comparison between our proposed LADiBI and the existing literature.
Method Family Method Prior Type Diverse Image Prior No Additional Training Flexible Operator
Pan-$\ell_{0}$ [28] Inductive bias × ×
Optimization-based Pan-DCP [29] Inductive bias × ×
Unsupervised SelfDeblur [30] Inductive bias × ×
MPRNet [44] Discriminative × ×
Supervised DeblurGANv2 [22] GAN × ×
BlindDPS [8] Pixel diffusion × × ×
GibbsDDRM [27] Pixel diffusion × ×
Diffusion-based LADiBI (Ours)
Text-to-image
latent diffusion
Diffusion for Inverse Problem Solving
Diffusion models, also known as score-based generative models [38, 16], generate clean data samples ${\bm{x}}_{0}$ by iteratively refining noisy samples ${\bm{x}}_{t}$ using a time-dependent score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$. This score function is usually parametrized as a noise predictor $\epsilon_{\theta}({\bm{x}}_{t},t)$ and can be used to produce the clean data samples by iteratively denoising the noisy samples from the noise prediction. Particularly, a popular sampling algorithm DDIM [35] adopts the update rule
${\bm{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{{\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{intermediate estimation of ${\bm{x}}_{0}$, denoted as }{\bm{x}}_{0|t}}$
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{x}}_{t},t)+\sigma_{t}\epsilon$ (1)
that consists of an intermediate estimation of the clean data in order to perform fast sampling.
Because of this iterative sampling process and its close connection to the score function, many efforts attempt to use unconditionally pretrained diffusion models for conditional generation [38, 26, 11, 37], especially inverse problem solving [9, 7, 19, 36], without additional training. Generally, when attempting to sample from the conditional distribution $p({\bm{x}}|{\bm{y}})$, they decompose its score as
$\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t}|{\bm{y}})=\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ (2)
Since $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be obtained from an unconditionally pre-trained diffusio model, these methods usually aim at deriving an accurate approximation for $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$.
As text-to-image latent diffusion models become increasingly popular and powerful  [33, 2, 45], recent work has explored using these large language-conditioned models for inverse problem-solving. In particular, MPGD [14] addresses noisy inverse problems by leveraging the natural manifold preserving property of the latent diffusion models. Specifically, it modifies the intermediate clean latent estimate ${\bm{z}}_{0|t}$ using the following posterior update rule
${\bm{z}}_{0|t}={\bm{z}}_{0|t}-c_{t}\nabla_{{\bm{z}}_{0|t}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ (3)
where ${\mathbf{D}}({\bm{z}}_{0|t})$ represents the decoded intermediate clean image estimation and $c_{t}$ is the step size hyperparameter to account for the unknown noise level $\sigma$. The $L_{2}$ loss $\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ is induced by the additive Gaussian noise assumption from conventional inverse problem setting. Although these diffusion-based methods performs well on diverse data distributions and tasks, they all require the operator ${\mathcal{A}}_{\hat{\phi}}$ to be known.
Blind Inverse Problem
Blind inverse problems aim to recover unknown data ${\bm{x}}\in\mathbb{R}^{d}$ from measurements ${\bm{y}}\in\mathbb{R}^{m}$, typically modeled as:
${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$ (4)
where ${\mathcal{A}}_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is the forward operator parameterized by unknown function $\phi$, and ${\bm{n}}\sim{\mathcal{N}}(0,\sigma^{2}I)\in\mathbb{R}^{m}$ represents measurement noise with variance $\sigma^{2}I$. Blind inverse problems are more difficult due to the joint estimation of ${\bm{x}}$ and ${\mathcal{A}}_{\phi}$, and are inherently ill-posed without further assumptions. As a result, existing methods typically incorporate different assumptions about the priors of the target image data as well as the unknown operator.
Conventional methods use hand-crafted functions as image and operator prior constraints [5, 29, 28, 20, 24]. They often obtain these functional constraints by observing certain characteristics (e.g. clear edges and sparsity) unique to certain distributions that are usually considered as the target data distribution (e.g. clear and well-calibrated photos of the real world) and the operator distribution (e.g. some blurring kernels). However, not only are these hand-made functions not generalizable, they also often require significant manual tuning efforts for each individual image.
With the rise of deep learning, neural network has become a popular choice for parameterizing priors [40, 12, 30, 22, 44, 21]. In particular, Kupyn et al. [21, 22], Zamir et al. [44] use supervised learning on data-measurement pairs to jointly learn the image-operator distributions, and Ulyanov et al. [40], Ren et al. [30] apply unsupervised learning solely on measurements by leveraging deep convolutional networks’ inherent image priors. These methods offer significant improvement over traditional approaches. However, their learning procedures require separate data collection and model training for each image distribution and task, making them resource-intensive. Although some attempt to generalize by training on a wide range of data and measurements, our experiments show that they still struggle to adapt to unseen image or operator distributions.
Recently, inspired by advances in diffusion models for inverse problem solving, numerous efforts have incorporated the generative priors from pre-trained diffusion models. However, most of these methods still lack generalizability [8, 27, 34, 23]. For instance, Chung et al. [8], Sanghvi et al. [34] require training separate operator priors, while Murata et al. [27] remains training-free but rely on the operator kernel’s SVD for feasible optimization. These approaches generally depend on linear assumptions about the operator and well-trained diffusion models tailored to specific image distributions, limiting their ability to generalize across diverse image and operator types.
Concurrent to our work, two methods [39, 1] attempt to address these limitations. However, Tu et al. [39] only slightly relaxes the operator constraints to linear function with additive masking. It also only uses pixel diffusion rather than more powerful latent diffusion models. Anonymous [1] relies on hand-crafted update rules for each task, lacking a truly generalizable framework. More importantly, neither method leverages the flexibility provided by text prompts in large text-to-image diffusion models.
Table 1 summarizes the conceptual difference between out method and popular approaches in current literature. By leveraging large-scale pre-trained text-to-image latent diffusion models and our new posterior sampling algorithm, our method offers the most generalizability across diverse image and operator distributions with no additional training.
3 Method
In this paper, we aim to tackle the problem of blind inverse problem solving defined in Equation 4. Our solution (Algorithm 1) has the following desiderata: (1) No additional training: our method should not require data collection or model training, (2) Adaptability to diverse image priors: the same model should apply to various image distributions, (3) Flexible operators: no assumptions about the operator’s functional form, such as linearity or task-specific update rules, should be necessary. To make this problem feasible, we assume access to open-sourced pre-trained models.
To tackle this problem, we first formulate it as a Bayesian inference problem where the optimal solution is to sample from the posterior
$p({\bm{x}},{\mathcal{A}}_{\phi}|y)\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi}).$ (5)
This formulation allows us to decompose this problem into two parts: first obtaining a reliable estimation of the prior $p({\mathcal{A}}_{\phi})$ and then maximizing the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$. This makes diffusion-based framework the ideal approach as its posterior sampling process naturally separates these two stages (by Equation 2).
Algorithm 1 Our algorithm LADiBI
1: /* Initialize latent with encoded measurement ${\bm{y}}$ SDEdit */
2:${\bm{z}}_{T_{s}}=\sqrt{\bar{\alpha}_{T_{s}}}{\mathbf{E}}(y)+\sqrt{1-\bar{\alpha}_{T_{s}}}\epsilon_{T_{s}}$ for $\epsilon_{T_{s}}\sim\mathcal{N}(0,I)$
3:Initialize $\hat{\phi}$ with a fixed operator prior or Algorithm 2
4:for $t=T_{s},\dots,1$ do
5:     /* Use time-traveling for more stable results */
6:    for $j=1,\dots,M$ do
7:          /* Use CFG to obtain accurate prior */
8:         Calculate $\hat{\epsilon}_{\theta}({\bm{z}}_{t},t)$ with Eq. 7
9:         ${\bm{z}}_{0|t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}_{\theta}({\bm{z}}_{t},t))$
10:         if $j\equiv 0\pmod{2}$ then
11:              /* Perform MPGD with the estimated ${\mathcal{A}}_{\hat{\phi}}$/
12:             Update ${\bm{z}}_{0|t}$ with Eq. 8
13:         end if
14:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t},t)+\sigma_{t}\epsilon_{t}$ for $\epsilon_{t}\sim\mathcal{N}(0,I)$
15:         ${\bm{z}}_{t}=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}{\bm{z}}_{t}+\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\epsilon^{\prime}_{t}$ for $\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I)$
16:    end for
17:     /* Periodically update $\hat{\phi}$ with ${\bm{z}}_{0|t}$ and Eq. 9 */
18:    if $t\equiv 0\pmod{5}$ then
19:         for $k=1,\dots,K$ do
20:             $\hat{\phi}=\text{Adam}(\hat{\phi},L_{\phi}({\bm{z}}_{0|t},\hat{\phi}))$
21:         end for
22:    end if
23:end for
24:return ${\bm{x}}_{0}={\mathbf{D}}({\bm{z}}_{0})$
3.1 Obtaining a better Prior
Most existing diffusion-based methods [8, 27, 34, 23, 1, 39] avoid modeling the joint prior and instead further decompose the goal distribution as $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}})p({\mathcal{A}}_{\phi})$, under the assumption that the target data distribution $p({\bm{x}})$ independently from the operator distribution $p({\mathcal{A}}_{\phi})$. In this way, they can separately model the image and operator distribution by pre-trained diffusion models or fixed distributions and thus bypass the combinatorial complexity. However, is this truly a fair assumption?
Consider the task of deblurring and super-resolution. A low-resolution blurry measurement may correspond to a high-resolution blurry image (e.g. an out of focus DSLR photo) if the operator is a downsampler, or to a low-resolution clear image if the operator is a blur kernel. Similarly, different target image distribution can imply different operators: for example, the estimated blur kernel should differ if the target is a realistic photo versus an Impressionist painting, even when both share the same measurement.
These observations suggest that $p({\bm{x}})$ and $p({\mathcal{A}}_{\phi})$ should be correlated, and previous methods can make this independence assumption because they implicitly restrict the types of operations and target images they handle. In other words, while this independence assumption reduces modeling complexity, it also significantly limits the flexibility of the resulting algorithms. Consequently, most existing methods require additional training or parametrization of the operator prior and need to switch the image prior models when the target distribution changes.
Therefore, to build a unified training-free algorithm, we should directly model $p({\bm{x}},{\mathcal{A}}_{\phi})$, or in the case of diffusion models, $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$. On the first glance, modeling $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ appears intractable. However, with large-scale pre-trained text-to-image diffusion models now widely available, we have a surprisingly simple solution:
In practice, the target data distribution (e.g. “a high-resolution photo”) and the forward operation (e.g. “Gaussian blur”) can often be effectively described in natural language. We can specify the target distribution by using its description as a positive prompt. Specifying the operator distribution is less direct, as these pre-trained models are designed to represent image distributions rather than operators. However, by leveraging classifier-free guidance (CFG) [15], we can implicitly define the operator distribution through negative prompts. For example, if the operator is Gaussian blur, an effective negative prompt can be “blurry image, low quality”. Formally, denote positive prompt as ${\bm{c}}_{+}$ and negative prompt as ${\bm{c}}_{-}$, we can approximate $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ as
$\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-})\\ +\gamma(\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{+})-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-}))$ (6)
where $\gamma>1$ is a weighting hyperparameter. This is equivalent to approximating $p({\bm{x}},{\mathcal{A}}_{\phi})$ as a distribution that is proportional to $p({\bm{x}}|{\bm{c}}_{-})^{1-\gamma}p({\bm{x}}|{\bm{c}}_{+})^{\gamma}$.
When using latent diffusion models parameterized by $\theta$, this corresponds to having the empirical noise prediction
$\small\hat{\epsilon}_{\theta}({\bm{z}}_{t},t)=\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-})+\gamma(\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{+})-\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-}))$ (7)
This way, our method not only models this seemingly extremely intractable joint distribution in a remarkably simple way, it also enable wide range of applications with diverse target image and operator distributions by leveraging the knowledge in large-scale pre-trained text-to-image diffusion models like Stable Diffusion[31].
3.2 Obtaining a better Posterior
Algorithm 2 General Operator Initialization
1:for $j=1,\dots,M$ do
2:     /* Initialize latent batch with encoded ${\bm{y}}$ SDEdit */
3:    $\epsilon^{(i)}_{T_{s}}\sim\mathcal{N}(0,I)$ for $i=1,2,\dots,N$
4:    ${\bm{z}}_{T_{s}}^{(i)}=\sqrt{\bar{\alpha}_{T_{s}}}{\mathbf{E}}(y)+\sqrt{1-\bar{\alpha}_{T_{s}}}\epsilon^{(i)}_{T_{s}}$
5:     /* Use MPGD to obtain latent estimations $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$/
6:    for $t=T_{j},\dots,1$ and all $i$ in parallel do
7:         Calculate $\hat{\epsilon}_{\theta}({\bm{z}}_{t}^{(i)},t)$ with Eq. 7
8:         ${\bm{z}}_{0|t}^{(i)}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\bm{z}}_{t}^{(i)}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}_{\theta}({\bm{z}}_{t}^{(i)},t))$
9:         if $j\neq 1$ then
10:             Update ${\bm{z}}^{(i)}_{0|t}$ with Eq. 8
11:         end if
12:         ${\bm{z}}^{(i)}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}^{(i)}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t}^{(i)},t)+\sigma_{t}\epsilon_{t}^{(i)}$ for $\epsilon^{(i)}_{t}\sim\mathcal{N}(0,I)$
13:    end for
14:     /* Update $\hat{\phi}$ with $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$ and Eq. 9 */
15:    for $k=1,\dots,K$ do
16:         $\hat{\phi}=\text{Adam}(\hat{\phi},\frac{1}{N}\sum_{i=1}^{N}L_{\phi}({\bm{z}}_{0}^{(i)},\hat{\phi}))$
17:    end for
18:end for
19:return $\hat{\phi}$
In addition to a strong prior, our final output image should also satisfy the measurement constraint.Given the problem setup in Equation 4, the measurements are subject to additive Gaussian noise ${\bm{n}}$, hence $\log p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})=-\frac{1}{2\sigma^{2}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\bm{x}})\|^{2}_{2}$. When ${\mathcal{A}}_{\phi}$ is known, we can use the MPGD update rule in Equation 3 to perform posterior sampling.
However, since ${\mathcal{A}}_{\phi}$ is not known, the true parameters $\phi$ is often approximated by another set of parameters $\hat{\phi}$. This approximation is usually addressed by one of the two strategies: an alternating optimization scheme that jointly approximates ${\bm{x}}$ and $\phi$, or obtaining a reliable $\hat{\phi}$ first then solving a non-blind inverse problem. The first approach is well-suited for training-free settings as it can be well integrated with diffusion-based methods, but it is often highly sensitive to $\hat{\phi}$ initialization and tuning. The second approach can perform well if a strong $\hat{\phi}$ is obtained, though it usually requires training and additional assumptions or restrictions.
We propose to combine the two strategies above: we first obtain a reliable initial $\hat{\phi}$, and then perform an alternating optimization to iteratively refine both the operator parameter and data estimations throughout the diffusion process.
$\hat{\phi}$ Initialization
In some cases, initializing the operator parameters is straightforward – for example, for a slight blur, an identity function can be a reasonable stating point. Alternatively, reliable operator priors, such as the pre-trained prior in BlindDPS [8] or the Gaussian prior in GibbsDDRM [27] for linear operators, can also provide effective initializations for simple kernels. However, when the operator requires complex functions like neural networks, obtaining good initializations becomes more challenging.
To address this, we propose a new algorithm for general operator initialization by drawing further inspiration from alternating optimization algorithms. Note that since the goal here is only to initialize $\hat{\phi}$, the quality of intermediate ${\bm{x}}$ estimates is less critical as long as they provide a useful signal. Unlike other diffusion-based methods that alternate optimization targets at each diffusion step, we use SDEdit [26] and MPGD [14] with very few diffusion steps to quickly obtain a batch of ${\bm{x}}$ estimates. We then perform maximum likelihood estimation (MLE) on these estimates to update $\hat{\phi}$ using Adam optimizer. As detailed in Algorithm 2, repeating this process can leverage the diffusion prior to quickly converge to a reliable starting point for $\hat{\phi}$.
Notice that we only assume that $\phi$ can be approximated by differentiable functions, allowing ${\mathcal{A}}_{\hat{\phi}}$ to be parameterized by general model families such as neural networks. Moreover, as mentioned earlier, any well-performing operator priors can be seamlessly integrated into our framework.
Iterative Refinement
After initializing $\hat{\phi}$, we then perform another alternating optimization scheme to refine both the operator and the data. In particular, we alternate between updating ${\bm{z}}_{0|t}$ using MPGD guidance with ${\mathcal{A}}_{\hat{\phi}}$ fixed and updating the MLE estimation of $\hat{\phi}$ using Adam with ${\bm{z}}_{0|t}$ fixed throughout the diffusion process.
Since unlike GibbsDDRM, which uses Langevin dynamics to update the operator, we solve for a local optimum. Therefore, it’s reasonable not to update the operator too frequently. Empirically, we find that periodic updates to the operator in combination of the time-traveling technique [25, 42] yield the best results
In addition, since MPGD supports any differentiable loss function, we can incorporate techniques like regularization to further improve the visual quality. In practice, we use
$\small{\bm{z}}_{0|t}={\bm{z}}_{0|t}-c_{t}\nabla_{{\bm{z}}_{0|t}}(\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}+\lambda_{\bm{z}}\text{LPIPS}({\mathbf{D}}({\bm{z}}_{0|t}),{\bm{y}}))$ (8)
as the MPGD update rule and
$L_{\phi}({\bm{z}}_{0|t},\hat{\phi})=\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}+\lambda_{\phi}\|\hat{\phi}\|_{1}$ (9)
as the Adam objective. Here $\text{LPIPS}(\cdot)$ denotes the LPIPS distance between two images and $\|\cdot\|_{1}$ denotes the $L_{1}$ regularization term. $\lambda_{\bm{z}}$ and $\lambda_{\phi}$ are adjustable hyperparameters.
By combining large-scale language-conditioned diffusion priors, effective operator initializations, and the iterative refinement process described above, we introduce LADiBI (Algorithm 1), a new training-free algorithm for blind inverse problem solving that supports diverse target image distributions and flexible operators.
4 Experiments
4.1 Experimental Setup
We empirically verify the performance of our method with two linear deblurring tasks, Gaussian deblurring and motion deblurring, and a non-linear restoration task, JPEG decompression. Here we provide the details of our experiments.
Datasets
We conduct quantitative comparisons on the datasets FFHQ 256 $\times$ 256 [18] and AFHQ-dog 256 $\times$ 256 [6]. We use a size of 1000 images for FFHQ and 500 for AFHQ, following the setting in Murata et al. [27].
Baselines
We compare our method against other state-of-the-art approaches as baselines. Specifically, we choose Pan-l0 [28] and Pan-DCP [29] as the optimization-based method, SelfDeblur [30] as the self-supervised approach, PRNet [44] and DeblurGANv2 [22] as the supervised baselines, and BlindDPS [8] and GibbsDDRM [27] as the diffusion-based methods. All baselines are experiments using their open-sourced code and pre-trained models, whose details are provided in the appendix.
Implementation Details
We conduct all experiments an NVIDIA A6000 GPU. We use Stable Diffusion 1.4 [31], DDIM $200$-step noise scheduling, and $T_{s}=150,M=4$ for all tasks. We set positive prompts as “a clear headshot of a person” and “a clear headshot of an animal” for FFHQ and AFHQ experiments respectively. We adjust negative prompts according to each individual task, and the details will be provided in the following sections.
Evaluation Metrics
We evaluate our method and the baselines with three quantitative metrics: LPIPS [46], PSNR and KID [4]. Each metric has its own significance and reveals a different set of characteristics about performance. In particular, PSNR directly measures the pixel-by-pixel accuracy of the reconstructed image compared to the original, while LPIPS measures how similar two images are in a way that reflects human perception. Finally, KID evaluates the image fidelity on a distribution level.
4.2 Blind Linear Deblurring
We first test our method against the baselines on linear deblurring, which is the design space of most baselines. We evaluate all methods on two types of blur kernels: Gaussian blur and motion blur.
Setup
To generate measurements, we apply randomly generated motion blur kernels with intensity 0.5 and Gaussian blur kernels with standard deviation 3.0 respectively. The final measurements are derived by applying a pixel-wise Gaussian noise with $\sigma=0.02$. We use a $61\times 61$ convolutional matrix as $\hat{\phi}$ and we initialize it as a Gaussian kernel of intensity 6.0 following Murata et al. [27]. We adjust the negative prompts to the potential negative image artifacts induced by the degradation operation, such as “lowres, blurry, DSLR effect, poorly drawn face, deformed, bad proportions, …” for Gaussian blur, “shaken image, morbid, mutilated, text in image, disfigured, …” for motion blur. Full prompts are included in the appendix.
Results
Table 2: Quantitative results on blind linear deblurring tasks. “GibbsDDRM” denotes the results from directly running their open-sourced code and “GibbsDDRM*” denotes the results after adjusting their code to match with the ground truth kernel padding.
FFHQ AFHQ
Method Motion Gaussian Motion Gaussian
[…]
SelfDeblur 0.732 9.05 0.1088 0.733 8.87 0.0890 0.742 9.04 0.0650 0.736 8.84 0.0352
MPRNet 0.292 22.42 0.0467 0.334 23.23 0.0511 0.324 22.09 0.0382 0.379 21.97 0.0461
DeblurGANv2 0.309 22.55 0.0411 0.325 26.61 0.0227 0.323 22.74 0.0350 0.340 27.12 0.0073
Pan-l0 0.389 14.10 0.1961 0.265 20.68 0.1012 0.414 14.16 0.1590 0.276 21.04 0.0320
Pan-DCP 0.325 17.64 0.1323 0.235 24.93 0.0490 0.371 17.63 0.1377 0.297 25.11 0.0263
BlindDPS 0.246 20.93 0.0153 0.216 25.96 0.0205 0.393 20.14 0.0913 0.330 24.79 0.0268
GibbsDDRM 0.293 20.52 0.0746 0.216 27.03 0.0430 0.303 19.44 0.0265 0.257 24.01 0.0040
LADiBI (Ours) 0.230 20.96 0.0084 0.197 21.08 0.0068 0.262 21.20 0.0132 0.204 24.33 0.0065
GibbsDDRM* 0.199 22.36 0.0309 0.155 27.65 0.0252 0.278 19.00 0.0180 0.224 21.62 0.0040
Figure 2: Qualitative results on blind linear deblurring tasks. From top to bottom we showcase examples from motion deblur on FFHQ, Gaussian deblur on FFHQ, motion deblur on AFHQ, and Gaussian deblur on AFHQ respectively.
Table 2 and Figure 2 show results on blind linear deblurring. Our method matches the performance the state-of-the-art method GibbsDDRM, which is explicitly designed to solving linear problems using SVDs. In fact, GibbsDDRM’s highly specialized design makes it so sensitive that even small discrepancies (e.g. kernel padding) between their modeling assumption and the ground truth operator can result in significant performance degradation. Moreover, although BlindDPS and GibbsDDRM use diffusion models that are trained on the exact data distributions tested, our method can still match or outperform them with the large-scale pre-trained model. In addition, even though supervised methods can obtain higher PSNR, our method produces the fewest artifacts, which is also validated by the LPIPS and KID scores. Overall, our method offers a consistently competitive performance on linear tasks, even though our method is designed for more general applications.
4.3 Blind JPEG Decompression
We further compare our method with baselines on JPEG decompression, a particularly challenging task due to the non-linear, non-differentiable nature of JPEG compression. Unlike traditional settings, our experiments provide no task-specific knowledge, such as the compression algorithm, quantization table, or factors – algorithms rely solely on the measurement images.
Setup
We generate measurements using JPEG compression with a quantization factor of 2. We use a neural network with a 3-layer U-net [32] to parametrize the operator. The operator is initialized using Algorithm 2 with $M=8$ and $N=4$. We use phrases including “pixelated, low quality, jpeg artifacts, deformed, bad proportions, …” as negative prompt, whose full version is provided in the appendix.
Results
Table 3: Quantitative results on blind JPEG decompression task.
Method FFHQ AFHQ
LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
Figure 3: Qualitative results on the blind JPEG decompression task.
Table 3 and Figure 3 present the results on the JPEG decompression task. It is evident through both quantitative and qualitative results that our method is the only one capable of enhancing the fidelity of the images and maintaining consistency to measurements. Unlike the baselines, which struggle with this task due to their limited posterior formulations, our flexible framework adapts to approximate this challenging operator and produces high quality images.
4.4 Ablation Studies
Table 4: Ablation study on our proposed algorithm.
Removed Component LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
Text prompts 0.500 18.88 0.0483
MPGD guidance 0.470 17.80 0.0434
$\hat{\phi}$ Initialization 0.355 19.74 0.0250
LADiBI 0.315 21.12 0.0216
To showcase the effectiveness of each part of our algorithm, we conduct ablation study on AFHQ dataset with JPEG decompression task. In particular, we test the importance of suitable ad-hoc prompts, the use of MPGD guidance, and the operator initialization for neural network ${\mathcal{A}}_{\hat{\phi}}$. in-between diffusion steps. We keep SDEdit in all settings to encode the measurement information for fair comparisons. As shown in Table 4, each of the aforementioned components plays a significant role in our scheme and is indispensable for producing high quality results.
4.5 Style Preserving Deblurring
Figure 4: Qualitative results on blind deblurring Monet paintings.
In order to demonstrate our method’s ability to solve blind inverse problems from a wider range of image distributions, we present indicative examples of Gaussian and motion deblurring on Monet paintings in Figure 4. We use the same negative prompts as in the previous experiments and “a portrait of a person as a Monet style painting” as positive prompts. We notice that, although the images portray human faces, the baseline method using models trained on realistic human face images cannot accurately solve the problem, while our method effectively generates images consistent with both the measurement image and the painting style present in the ground truth image.
5 Conclusion
In this work, we propose LADiBI, a new training-free algorithm to solving blind inverse problems in image restoration using a large-scale pre-trained text-to-image diffusion model. Our method leverages text prompts as well as posterior guidance on intermediate diffusion steps to guide the generation process towards clean reconstructions based on degraded measurements, in cases where the degradation operation is unknown. Experimental results demonstrate that LADiBI achieves competitive performance on tasks with diverse operator and is also capable of covering a wider range of image distributions compared to baselines.
Acknowledgment
This work is sponsored by Sony AI and is also supported in part by the ONR grant N000142312368.
Appendix A Additional Implementation Details
We offer more detailed information regarding the implementation setup for Algorithms 1 and 2. All experiments are conducted using an NVIDIA A6000 GPU. In an effort to make our methodology generalizable and extensible to other inverse problems, we attempt to maintain the same hyper-parameter values across tasks wherever possible. Each hyper-parameter value has been chosen after conducting preliminary experiments for a specific range and opting for the value that offers the best performance. Indicative examples of such experiments are shown in C. In particular, we run Algorithm 1 for $T_{s}=150$ timesteps while performing 4 repetitions as part of the time-traveling strategy. We encode the measurement as ${\bm{x}}_{T_{s}}$ by applying the forward diffusion process up to timestep 800. In parallel with the reverse diffusion process, we update $\hat{\phi}$ every 5 timesteps, each time conducting $K=150$ gradient steps. We adust the $c_{t}$ and the $\lambda_{\phi}$ parameters of Equations 8 and 9 to 30 and 2 respectively. In terms of the operator initialization process described by Algorithm 2, we make use of a batch of $N=4$ samples and run $M=8$ iterations, each comprising $T_{j}=60$ timesteps.
Figure 5: A schematic overview of Algorithm 1.
Figure 6: A schematic overview of Algorithm 2.
Additionally, schematic overviews of Algorithms 1 and 2 are presented in Figures 5 and 6.
Furthermore, there are some parameters in our approach for which employing a task-aware setup strategy is essential for state-of-the-art performance. These are the prompt sentences as well as the architecture of $\hat{{\phi}}$. Here we provide details with respect to these parameters according to each restoration task:
Motion Deblurring
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “shaken image, motion blur, pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
  • $\hat{\phi}$ architecture: A single $61\times 61$ convolutional block with 3 input and 3 output channels.
Gaussian Deblurring
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “blurry, gaussian blur, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
  • $\hat{\phi}$ architecture: A single $61\times 61$ convolutional block with 3 input and 3 output channels.
JPEG Decompression
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, jpeg artifacts, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
  • $\hat{\phi}$ architecture: A neural network with a typical 3-layer U-net [32] architecture. Each layer consists of 2 convolutional blocks of size $3\times 3$ with ReLU activations and number of input and output channels ranging from 32 to 128.
Appendix B Additional Results
Table 5: Additional quantitative results on blind JPEG decompression task.
Method FFHQ AFHQ
LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
Pan-l0 0.787 12.72 0.2190 0.825 13.15 0.2168
Pan-DCP 0.710 14.91 0.2318 0.673 18.27 0.2472
SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
Figure 7: Additional qualitative results on motion deblurring task.
Figure 8: Additional qualitative results on Gaussian deblurring task.
Figure 9: Additional qualitative results on JPEG decompression task.
Figure 10: Additional qualitative results on painting restoration.
We present more detailed experimental results on all benchmark tasks. In specific, Table 5 shows quantitative results on the JPEG decompression task, including the optimization-based Pan-l0 and Pan-DCP methods. Moreover, Figures 7, 8 and 9 offer more qualitative comparisons against baseline methods on motion deblurring, gaussian deblurring and JPEG decompression respectively. Finally, in Fig. 10 we include comparisons against selected baseline methods on the 3 benchmark tasks using a set of Monet and Van Gogh portrait paintings as ground truth images.
Appendix C Additional Ablation Study
Figure 11: Ablation study on diffusion timestep for encoding the measurement.
Figure 12: Ablation study on number of time-traveling steps.
Figure 13: Comparison between linear and deep operator architecture on linear tasks.
We carefully adjust each parameter of our methodology to the value that offers the best overall results by conducting comparative experiments and in this section we offer some indicative examples of such tests. All hyper-parameter tuning experiments have been performed on the motion deblurring task using the AFHQ dataset.
Figure 11 presents both the LPIPS and KID score for the value of the timestep at which we encode the measurement using the forward diffusion process. Both metrics form a U-shape curve, which results from the following fact about encoding the measurement image; if we make the encoding timestep too large, most of the information about the measurement has been replaced by noise which does not allow the sampling process to leverage useful features of the degraded image. On the other hand, if the timestep is too small, the sampling process is not equipped with enough scheduled variance to reach the target prior distribution.
In addition, Figure 12 presents the effectiveness of taking advantage of the time-traveling strategy. More time-traveling boosts the overall performance up to a specific value, after which we begin to notice a trade-off between perceptual clarity of the image and fidelity to the target distribution.
With respect to ablation studies, we are interested in evaluating the performance of our algorithm when using a task-agnostic implementation for ${\mathcal{A}}_{\hat{\phi}}$. Hence, we tested the neural network architecture employed for non-linear tasks at linear inverse problems and we showcase qualitative results in Figure 13. We observe that, although performing worse than LADiBI employed with a linear operator architecture, the task-agnostic method is capable of producing estimates of decent quality while also preserving a design that allows applicability to any image restoration task.
Appendix D Limitations and Future Works
In this section, we discuss the limitation of LADiBI and potential future works to address these problems.
Similar to many training-free diffusion posterior sampling algorithms [8, 27, 9, 14], our method is also sensitive to hyper-parameter tuning. We have provided details about our hyperparameter choices in the previous sections, and we will release our code in a public repository upon publication of this paper.
Another notable drawback of our method is that LADiBI requires significantly longer inference time in comparison to the best performing baseline (i.e. GibbsDDRM [27]) in order to obtain high quality restorations. With the general operator initialization, our algorithm can take around 5 minutes to complete on a single NVIDIA A6000 GPU while GibbsDDRM only takes around 30 seconds. Although this is justifiable by the larger optimization space that we operate on, investigating on how to reduce the inference time requirement is an interesting and critical next step for our work.
Lastly we would like to note that, while neural networks, as demonstrated in the previous sections, can serve as a general model family for various operator functional classes and achieve satisfactory results, obtaining state-of-the-art performance for linear tasks within a reasonable inference time still requires resorting to a linear kernel as the estimated operator. Exploring neural network architectures that can easily generalize across different operator functional classes while achieving state-of-the-art results efficiently remains an exciting direction for future work.
Appendix E Broader Impact Statement
Lastly, since our algorithm leverages large pre-trained image generative models, we would like to address the ethical concerns, societal impact as well as the potential harm that can be caused by our method when being used inappropriately.
As a consequence of using large pre-trained text-to-image generative models, our method also inherits potential risks associated with these pre-trained models, including the propagation of biases, copyright infringement and the possibility of generating harmful content. We recognize the significance of these ethical challenges and are dedicated to responsible AI research practices that prevent reinforcing these ethical considerations. Upon the release our code, we are committed to implement and actively update safeguards in our public repository to ensure safer and more ethical content generation practices.

### v2

@@
 Abstract
-Blind inverse problems, where both the target data and forward operator are unknown, are crucial to many computer vision applications.
-Existing methods often depend on restrictive assumptions such as additional training, operator linearity, or narrow image distributions, thus limiting their generalizability.
-In this work, we present LADiBI, a training-free framework that uses large-scale text-to-image diffusion models to solve blind inverse problems with minimal assumptions.
-By leveraging natural language prompts, LADiBI jointly models priors for both the target image and operator, allowing for flexible adaptation across a variety of tasks.
-Additionally, we propose a novel posterior sampling approach that combines effective operator initialization with iterative refinement, enabling LADiBI to operate without predefined operator forms.
-Our experiments show that LADiBI is capable of solving a broad range of image restoration tasks, including both linear and nonlinear problems, on diverse target image distributions.
-Figure 1: Our proposed LADiBI is a training-free blind inverse problem solving algorithm using large pre-trained text-to-image diffusion models.
+This paper considers blind inverse image restoration, the task of predicting a target image from a degraded source when the degradation (i.e.
+the forward operator) is unknown.
+Existing solutions typically rely on restrictive assumptions such as operator linearity, curated training data or narrow image distributions limiting their practicality.
+We introduce LADiBI, a training-free method leveraging large-scale text-to-image diffusion to solve diverse blind inverse problems with minimal assumptions.
+Within a Bayesian framework, LADiBI uses text prompts to jointly encode priors for both target images and operators, unlocking unprecedented flexibility compared to existing methods.
+Additionally, we propose a novel diffusion posterior sampling algorithm that combines strategic operator initialization with iterative refinement of image and operator parameters, eliminating the need for highly constrained operator forms.
+Experiments show that LADiBI effectively handles both linear and challenging nonlinear image restoration problems across various image distributions, all without task-specific assumptions or retraining.
+1 Introduction
+Figure 1: Our proposed LADiBI is a training-free blind inverse problem solving algorithm for image restoration using large pre-trained text-to-image diffusion models.
 LADiBI is applicable to a wide variety of image distribution as well as operators with minimal modeling assumptions imposed.
-1 Introduction
-Inverse problems are fundamental to many computer vision tasks, where the goal is to recover unknown data ${\bm{x}}$ from observed measurements ${\bm{y}}$.
-Formally, these problems can be expressed as ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$, where ${\mathcal{A}}$ is a operator representing the forward process parametrized by $\phi$, and ${\bm{n}}$ is the measurement noise.
-These problems are critical in fields such as medical imaging, computational photography and signal processing, as they are inherently tied to real-world scenarios like image decompression, deblurring, and super-resolution  [41, 43, 13, 17].
-Recently, deep learning-based methods  [30, 22, 44, 8, 7, 10] have emerged as powerful tools for solving inverse problems.
-In particular, diffusion models offer an elegant solution that incorporates pre-trained models in diverse conditional settings, providing flexible and scalable frameworks for reconstructing high-quality data by guiding the generation process through the measurement constraints [19, 36, 3, 37, 42, 14].
-However, most of these efforts focus on inverse problems where the operator ${\mathcal{A}}_{\phi}$ is assumed to be known.
-In practice, the forward operator is often unknown, leading to what are termed blind inverse problems.
-Solving these blind inverse problems remains a significant challenge due to their ill-posed nature.
-Consequently, current methods address this problem by introducing various assumptions about the prior distributions of the target data and unknown operator.
-These assumptions are typically incorporated through a combination of the following strategies: (1) using explicit formulas or specific neural network architectures to hand-craft inductive biases, (2) simplifying the operator with linear assumptions, (3) restricting target distributions to narrow image classes, (4) collecting measurement, operator or image datasets to train models tailored to specific tasks.
-While effective in certain cases, these solutions often entail significant computational and manual overhead, limited flexibility, and substantial barriers to practical deployment due to complex model training, costly data curation, unrealistic assumptions, and laborious hyperparameter tuning.
-These limitations raise a critical question: can we design an algorithm that applies to diverse operators and target image distributions without additional training or data collection?
-To tackle this challenge, we formulate this problem as a Bayesian inference problem, where the optimal solution is to sample from the posterior $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi})$.
-State-of-the-art diffusion-based methods often simplify this problem by assuming independence between the target image and operator distributions [8, 27, 23, 34].
-Although this independence assumption reduces the modeling difficulty, we challenge its validity by observing the correlation between the target image distribution and the operator distribution, and argue that this assumption heavily restricted the flexibility of the current diffusion-based approaches.
-To address this problem effectively, we first propose to estimate the joint distribution $p({\bm{x}},{\mathcal{A}}_{\phi})$ instead.
-While this estimation seems intractable on the first glance, there actually exists a surprisingly simple solution – large-scale pre-trained text-to-image diffusion models already capture diverse target data and measurement distributions.
-Moreover, many inverse problems can be described using natural language, allowing tasks like deblurring or super-resolution to be expressed through prompts such as “high-definition, clear image” for the target and “blurry, low-quality” for the measurement.
-By simply prompting with classifier-free guidance [15], we can approximate the score of this challenging joint prior distributions across a wide range of image and operator types using the same base model, significantly extending the flexibility of the current diffusion-based framework.
-In addition to a strong prior, an effective posterior sampling technique is also crucial for meeting measurement constraints.
-Existing methods generally fall into two categories: Gibbs or EM-style alternating optimization [27, 23], which suits diffusion-based setups without additional training but is sensitive to operator initialization and tuning, and a two-step approach that first estimates the operator followed by solving the non-blind problem [34], which often requires training and strong assumptions.
-To obtain the best of both worlds, we propose a hybrid approach: first, we obtain a reasonable initial estimate of the operator parameters, and then apply the alternating optimization to iteratively refine both operator parameters and data estimates throughout the diffusion process.
-Obtaining a reliable initial operator estimate can be difficult for broad functional classes like neural networks.
-To address this, we introduce a novel initialization scheme that leverages the pseudo-supervision signals from multiple lower-quality data estimation generated by fast posterior diffusion sampling.
-This approach eliminates the assumption of a specific form of operators, allowing for nonlinear blind inverse problem solving and flexible operator parametrization.
-Combining the strong prior with the effective posterior sampling technique, we introduce Language-Assisted Diffusion for Blind Inverse problems (LADiBI), a training-free method that leverages large-scale text-to-image diffusion models to solve a broad range of blind inverse problems with minimal assumptions.
-LADiBI can be directly applied across diverse data distributions and allows for easy specification of task-specific modeling assumptions through simple prompt adjustments.
-Algorithm 1 provides an overview of LADiBI, which can be easily adapted from the standard inference algorithm used in popular text-to-image diffusion models.
-Unlike existing methods, LADiBI requires no model retraining or reselection for different target data distributions or operator functions; instead, all prior parameterization is encoded directly in the prompt, which users can adjust as needed.
+Image restoration is a critical problem in many fields such as medical imaging and computational photography, as it addresses real-world challenges including image decompression, deblurring, and super-resolution  (Yuan et al., 2007; Greenspan, 2009; Isaac and Kulkarni, 2015).
+These restoration tasks can be formulated as inverse problems, where the goal is to recover unknown image data $\bm{x}$} from observed measurements $\bm{y}$}.
+Formally, these problems can be expressed as ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$, where $\mathcal{A}$} is an operator representing the forward degradation process parametrized by $\phi$, and $\bm{n}$} is the measurement noise.
+Being widely applicable, this problem has attracted numerous solutions, ranging from methods with handcrafted inductive biases to deep learning, especially diffusion-based techniques (Kawar et al., 2022; Song et al., 2022; Bansal et al., 2023; Yu et al., 2023; He et al., 2024).
+However, most existing research focuses on the settings where the operator ${\mathcal{A}}_{\phi}$ is known.
+In practice, the operator is often unknown, leading to what are termed blind inverse problems that present significant challenges due to their ill-posed nature.
+Current methods attempt to address this problem through several restrictive strategies: (1) introducing hand-crafted inductive biases via explicit formulas or specialized neural architectures (Pan et al., 2018, 2017; Ren et al., 2020), (2) simplifying operators with linear assumptions (Chung et al., 2023a; Murata et al., 2023), (3) constraining target distributions to narrow image classes (Chihaoui et al., 2024; Laroche et al., 2024), or (4) training task-specific models on curated measurement, operator or image datasets (Zamir et al., 2021; Kupyn et al., 2019).
+While effective in specific scenarios, these approaches suffer from limited flexibility and significant deployment barriers due to costly training and laborious hyperparameter tuning.
+This raises a fundamental question: can we develop a more generalizable algorithm capable of handling diverse degradation operators and image distributions without additional training or data collection?
+To tackle this challenge, we reframe the problem via Bayesian inference, where we sample from the posterior $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi})$.
+This formulation naturally decomposes the seemingly intractable problem into two simpler sub-problems: estimating the prior $p({\bm{x}},{\mathcal{A}}_{\phi})$ and sampling according to the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$.
+This decomposition is particularly well-suited for diffusion-based frameworks, where we can leverage pre-trained models as the prior and guide the sampling process via measurement constraints.
+However, while pre-trained models typically exist for image, adequate priors or training data for the operators are generally unavailable.
+Thus, existing methods further factorize the joint prior into independent distributions and impose simplifying assumptions such as linearity on the operator, which severely restrict their generality and flexibility.
+Instead, our approach is motivated by the key observation that, common restoration tasks can be intuitively described through natural language (e.g., “high-definition, clear image” for targets, “blurry, low-quality” for measurements).
+Moreover, large pre-trained text-to-image diffusion models already encapsulate rich distributions of both targets and measurements.
+Although this can, in some way, seem to render the approach less of a true “blind” solution, in practice we note that all methods for blind inverse problems require some assumptions over the space of transformations, and using English text to encode the joint is an extremely flexible and easy-to-use mechanism.
+Based on these insights, we propose a simple yet powerful method: using classifier-free guidance (Ho and Salimans, 2021), we can approximate the joint prior’s score across diverse images and operators using a single pre-trained text-to-image diffusion model, dramatically broadening the flexibility of existing frameworks.
+Our approach is particularly advantageous in blind settings, since the appropriate prompts can often be easily inferred directly from the measurements, and textual descriptions enable us to co-encode unknown degradations and desired outputs without training or handcrafted priors.
+In addition to estimating the prior score, effectively sampling from the posterior distribution is also crucial for ensuring that restored images satisfy measurement constraints.
+In blind inverse problems, achieving this requires accurately estimating both data and operator parameters, ideally with generalized parameter classes such as neural networks to improve flexibility.
+However, reliably initializing these highly unconstrained operator parameters can be challenging.
+To address this, we propose a novel co-optimization diffusion posterior sampling algorithm specifically tailored for blind inverse problems.
+Our method begins with a new initialization scheme that leverages pseudo-supervision signals derived from multiple lower-quality target image approximations generated by fast posterior diffusion sampling.
+We then iteratively refine both operator parameters and data estimates through an alternating optimization procedure integrated within the diffusion sampling process.
+This strategy eliminates restrictive assumptions about operator forms, thereby enabling nonlinear blind inverse problem solving with highly flexible operator parametrizations.
+Combining the text-conditioned prior with our effective posterior sampling, we introduce Language-Assisted Diffusion for Blind Inverse problems (LADiBI), a training-free method that leverages large-scale text-to-image diffusion models to solve a broad range of blind image restoration problems with minimal assumptions.
+LADiBI can be directly applied across diverse data distributions and allows for easy specification of task-specific assumptions through simple prompting.
+Algorithm 1 and Figure 2 provide an overview of LADiBI, which can be easily adapted from the standard inference algorithm used in popular text-to-image diffusion models.
+Unlike existing methods, LADiBI requires no model retraining or reselection for different data distributions or operator functions.
+Instead, all prior parameterization is encoded directly in the prompt, which users can adjust as needed.
 Notably, we do not assume linearity of the operator, making LADiBI, to the best of our knowledge, the most generalizable approach to blind inverse problem solving in image restoration.
-We evaluate LADiBI against state-of-the-art baselines across several blind image restoration tasks, including two linear inverse problems (motion deblurring and Gaussian deblurring) and a nonlinear inverse problem (JPEG decompression), on various image distributions, as illustrated in Figure 1.
-In the linear cases, our method matches the performance of the state-of-the-art approaches, which require extensive assumptions and model specifications.
-Additionally, LADiBI is the only method tested that successfully performs JPEG decompression without any information about the task, such as the compression algorithm, the quantization table or factors, relying solely on observations of the compressed images.
+We evaluate LADiBI against state-of-the-art baselines on a range blind image restoration tasks, including linear problem (e.g.
+motion and Gaussian deblurring) and nonlinear problem(e.g.
+JPEG decompression), across various image distributions, as illustrated in Figure 1.
+In the linear setting, our method matches the performance of the state-of-the-art approaches while requiring significantly fewer assumptions.
+In the nonlinear setting, LADiBI is the only method tested that can successfully perform JPEG decompression without any prior information of the task (such as the compression algorithm, quantization table or quantization factors), relying solely on observations of the compressed images.
 2 Background & Related Works
 Table 1: A conceptual comparison between our proposed LADiBI and the existing literature.
-Method Family Method Prior Type Diverse Image Prior No Additional Training Flexible Operator
-Pan-$\ell_{0}$ [28] Inductive bias × ×
-Optimization-based Pan-DCP [29] Inductive bias × ×
-Unsupervised SelfDeblur [30] Inductive bias × ×
-MPRNet [44] Discriminative × ×
-Supervised DeblurGANv2 [22] GAN × ×
-BlindDPS [8] Pixel diffusion × × ×
-GibbsDDRM [27] Pixel diffusion × ×
+Method Family Method Prior Type Diverse Image Prior Training-free Flexible Operator
+Pan-$\ell_{0}$ (Pan et al., 2017) Inductive bias × ×
+Optimization-based Pan-DCP (Pan et al., 2018) Inductive bias × ×
+Self-supervised SelfDeblur (Ren et al., 2020) Inductive bias × ×
+MPRNet (Zamir et al., 2021) Discriminative × ×
+Supervised DeblurGANv2 (Kupyn et al., 2019) GAN × ×
+BlindDPS (Chung et al., 2023a) Pixel diffusion × × ×
+BIRD (Chihaoui et al., 2024) Pixel diffusion × ×
+GibbsDDRM (Murata et al., 2023) Pixel diffusion × ×
 Diffusion-based LADiBI (Ours)
@@
 Diffusion for Inverse Problem Solving
-Diffusion models, also known as score-based generative models [38, 16], generate clean data samples ${\bm{x}}_{0}$ by iteratively refining noisy samples ${\bm{x}}_{t}$ using a time-dependent score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$.
-This score function is usually parametrized as a noise predictor $\epsilon_{\theta}({\bm{x}}_{t},t)$ and can be used to produce the clean data samples by iteratively denoising the noisy samples from the noise prediction.
-Particularly, a popular sampling algorithm DDIM [35] adopts the update rule
-${\bm{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{{\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{intermediate estimation of ${\bm{x}}_{0}$, denoted as }{\bm{x}}_{0|t}}$
-$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{x}}_{t},t)+\sigma_{t}\epsilon$ (1)
+Diffusion models, also known as score-based generative models (Song et al., 2021b; Ho et al., 2020), generate clean data samples ${\bm{x}}_{0}$ by iteratively refining noisy samples ${\bm{x}}_{t}$ using a time-dependent score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$.
+This score function is usually parametrized as a noise predictor $\epsilon_{\theta}({\bm{x}}_{t},t)$ and can be used to produce the clean data samples through iteratively denoising.
+Particularly, a popular sampling algorithm DDIM (Song et al., 2021a) adopts the update rule
+${\bm{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{{\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}({\bm{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{intermediate estimation of ${\bm{x}}_{0}$, denoted as }{\bm{x}}_{0|t}}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{x}}_{t},t)+\sigma_{t}\epsilon$ (1)
 that consists of an intermediate estimation of the clean data in order to perform fast sampling.
-Because of this iterative sampling process and its close connection to the score function, many efforts attempt to use unconditionally pretrained diffusion models for conditional generation [38, 26, 11, 37], especially inverse problem solving [9, 7, 19, 36], without additional training.
-Generally, when attempting to sample from the conditional distribution $p({\bm{x}}|{\bm{y}})$, they decompose its score as
+Many efforts attempt to use unconditionally pretrained diffusion for conditional generation (Song et al., 2021b; Meng et al., 2022; Dhariwal and Nichol, 2021), especially inverse problem solving (Chung et al., 2023b; Kawar et al., 2022; Song et al., 2022), without additional training.
+Generally, when sampling from $p({\bm{x}}|{\bm{y}})$, they decompose its score as
 $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t}|{\bm{y}})=\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ (2)
-Since $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be obtained from an unconditionally pre-trained diffusio model, these methods usually aim at deriving an accurate approximation for $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$.
-As text-to-image latent diffusion models become increasingly popular and powerful  [33, 2, 45], recent work has explored using these large language-conditioned models for inverse problem-solving.
-In particular, MPGD [14] addresses noisy inverse problems by leveraging the natural manifold preserving property of the latent diffusion models.
-Specifically, it modifies the intermediate clean latent estimate ${\bm{z}}_{0|t}$ using the following posterior update rule
+Since $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be obtained from an unconditionally pre-trained diffusion model, these methods usually aim at deriving an accurate approximation for $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$.
+Recent work has also explored text-to-image latent diffusion models for inverse problem-solving  (Saharia et al., 2022; Balaji et al., 2023; Zhang et al., 2024).
+In particular, MPGD (He et al., 2024) addresses inverse problems by leveraging the manifold preserving property of the latent diffusion models.
+Specifically, it modifies the intermediate clean latent estimate ${\bm{z}}_{0|t}$ with
 ${\bm{z}}_{0|t}={\bm{z}}_{0|t}-c_{t}\nabla_{{\bm{z}}_{0|t}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ (3)
-where ${\mathbf{D}}({\bm{z}}_{0|t})$ represents the decoded intermediate clean image estimation and $c_{t}$ is the step size hyperparameter to account for the unknown noise level $\sigma$.
-The $L_{2}$ loss $\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ is induced by the additive Gaussian noise assumption from conventional inverse problem setting.
-Although these diffusion-based methods performs well on diverse data distributions and tasks, they all require the operator ${\mathcal{A}}_{\hat{\phi}}$ to be known.
+where ${\mathbf{D}}({\bm{z}}_{0|t})$ represents the decoded intermediate clean image estimation and $c_{t}$ is the step size hyperparameter.
+The $L_{2}$ loss $\|{\bm{y}}-{\mathcal{A}}_{\phi}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}$ is induced by the additive Gaussian noise assumption from conventional inverse problem setting.
+Although these diffusion-based methods perform well on diverse data distributions and tasks, they all require the operator ${\mathcal{A}}_{\phi}$ to be known.
 Blind Inverse Problem
@@
 ${\bm{y}}={\mathcal{A}}_{\phi}({\bm{x}})+{\bm{n}}$ (4)
-where ${\mathcal{A}}_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is the forward operator parameterized by unknown function $\phi$, and ${\bm{n}}\sim{\mathcal{N}}(0,\sigma^{2}I)\in\mathbb{R}^{m}$ represents measurement noise with variance $\sigma^{2}I$.
+where ${\mathcal{A}}_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is the forward degradation operator parameterized by unknown function $\phi$, and ${\bm{n}}\sim{\mathcal{N}}(0,\sigma^{2}I)\in\mathbb{R}^{m}$ represents measurement noise with variance $\sigma^{2}I$.
 Blind inverse problems are more difficult due to the joint estimation of $\bm{x}$} and ${\mathcal{A}}_{\phi}$, and are inherently ill-posed without further assumptions.
 As a result, existing methods typically incorporate different assumptions about the priors of the target image data as well as the unknown operator.
-Conventional methods use hand-crafted functions as image and operator prior constraints [5, 29, 28, 20, 24].
+Conventional methods use hand-crafted functions as image and operator prior constraints (Chan and Wong, 1998; Pan et al., 2018, 2017; Krishnan et al., 2011; Levin et al., 2009).
 They often obtain these functional constraints by observing certain characteristics (e.g.
-clear edges and sparsity) unique to certain distributions that are usually considered as the target data distribution (e.g.
-clear and well-calibrated photos of the real world) and the operator distribution (e.g.
-some blurring kernels).
-However, not only are these hand-made functions not generalizable, they also often require significant manual tuning efforts for each individual image.
-With the rise of deep learning, neural network has become a popular choice for parameterizing priors [40, 12, 30, 22, 44, 21].
+clear edges and sparsity) unique to distributions that are usually considered as the target image (e.g.
+high definition natural images) and the operator (e.g.
+blurring kernels).
+However, not only are these hand-made functions not generalizable, they also often require significant manual tuning for each individual image.
+With the rise of deep learning, neural network has become a popular choice for parameterizing priors (Ulyanov et al., 2018; Gandelsman et al., 2019; Ren et al., 2020; Kupyn et al., 2019; Zamir et al., 2021).
 In particular, Kupyn et al.
-[21, 22], Zamir et al.
-[44] use supervised learning on data-measurement pairs to jointly learn the image-operator distributions, and Ulyanov et al.
-[40], Ren et al.
-[30] apply unsupervised learning solely on measurements by leveraging deep convolutional networks’ inherent image priors.
+(2019); Zamir et al.
+(2021) use supervised learning on data-measurement pairs to jointly learn the image-operator distributions, and Ulyanov et al.
+(2018); Ren et al.
+(2020) apply self-supervised learning solely on measurements using deep convolutional networks’ inherent image priors.
 These methods offer significant improvement over traditional approaches.
@@
 Although some attempt to generalize by training on a wide range of data and measurements, our experiments show that they still struggle to adapt to unseen image or operator distributions.
-Recently, inspired by advances in diffusion models for inverse problem solving, numerous efforts have incorporated the generative priors from pre-trained diffusion models.
-However, most of these methods still lack generalizability [8, 27, 34, 23].
+Recently, inspired by advances in diffusion models for inverse problem solving, numerous efforts have incorporated the priors from pre-trained diffusion models.
+However, most of these methods still lack generalizability (Chung et al., 2023a; Murata et al., 2023; Sanghvi et al., 2023; Laroche et al., 2024; Tu et al., 2024; WeiminBai et al., 2025).
 For instance, Chung et al.
-[8], Sanghvi et al.
-[34] require training separate operator priors, while Murata et al.
-[27] remains training-free but rely on the operator kernel’s SVD for feasible optimization.
+(2023a); Sanghvi et al.
+(2023) require training separate operator priors, while Murata et al.
+(2023) remains training-free but rely on the operator kernel’s SVD for feasible optimization.
 These approaches generally depend on linear assumptions about the operator and well-trained diffusion models tailored to specific image distributions, limiting their ability to generalize across diverse image and operator types.
-Concurrent to our work, two methods [39, 1] attempt to address these limitations.
-However, Tu et al.
-[39] only slightly relaxes the operator constraints to linear function with additive masking.
-It also only uses pixel diffusion rather than more powerful latent diffusion models.
-Anonymous [1] relies on hand-crafted update rules for each task, lacking a truly generalizable framework.
-More importantly, neither method leverages the flexibility provided by text prompts in large text-to-image diffusion models.
-Table 1 summarizes the conceptual difference between out method and popular approaches in current literature.
+Table 1 summarizes the conceptual difference between our method and popular approaches in current literature.
 By leveraging large-scale pre-trained text-to-image latent diffusion models and our new posterior sampling algorithm, our method offers the most generalizability across diverse image and operator distributions with no additional training.
@@
 In this paper, we aim to tackle the problem of blind inverse problem solving defined in Equation 4.
-Our solution (Algorithm 1) has the following desiderata: (1) No additional training: our method should not require data collection or model training, (2) Adaptability to diverse image priors: the same model should apply to various image distributions, (3) Flexible operators: no assumptions about the operator’s functional form, such as linearity or task-specific update rules, should be necessary.
+Our solution has the following desiderata: (1) No additional training: it should not require data collection or model training, (2) Adaptability to diverse image priors: the same model should apply to various image distributions, (3) Flexible operators: no assumptions about the operator’s functional form, such as linearity or task-specific update rules, should be necessary.
 To make this problem feasible, we assume access to open-sourced pre-trained models.
@@
 $p({\bm{x}},{\mathcal{A}}_{\phi}|y)\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}},{\mathcal{A}}_{\phi}).$ (5)
-This formulation allows us to decompose this problem into two parts: first obtaining a reliable estimation of the prior $p({\mathcal{A}}_{\phi})$ and then maximizing the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$.
-This makes diffusion-based framework the ideal approach as its posterior sampling process naturally separates these two stages (by Equation 2).
+This formulation allows us to decompose this problem into two parts: obtaining sampling to maximizes the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$.
+This makes diffusion-based framework the ideal approach as its posterior sampling process naturally separates these two stages.
+Figure 2: A schematic overview of LADiBI (Algorithm 1).
+3.1 Obtaining the prior score
+As established in Equation 2, diffusion-based approaches to inverse problems require access to the score of the prior distribution.
+For blind inverse problems, this extends to estimating the score of the joint distribution $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$, which presents significant challenges beyond standard image priors.
+While strong pre-trained diffusion models usually exist for target image distributions $p({\bm{x}})$, appropriate models or even training data for degradation operators ${\mathcal{A}}_{\phi}$ are largely unavailable, making estimations of the joint distribution $p({\bm{x}},{\mathcal{A}}_{\phi})$ particularly difficult.
+Therefore, most existing diffusion-based methods further decompose the joint prior into independent distributions $p({\bm{x}},{\mathcal{A}}_{\phi})\approx p({\bm{x}})p({\mathcal{A}}_{\phi})$ and impose restrictive assumptions, such as linearity constraints, on the operator (Chung et al., 2023a; Murata et al., 2023; Sanghvi et al., 2023; Laroche et al., 2024; WeiminBai et al., 2025; Tu et al., 2024).
+To estimate the operator distribution, they either rely on simulated training data tailored to specific operator forms or perform constrained optimization under fixed operator distributions.
+In addition, these methods typically require different pre-trained diffusion models for different image domains.
+For example, deblurring human faces versus animal faces necessitate different prior models.
+While effective in constrained settings, their reliance on domain-specific models and strong operator assumptions fundamentally limits their generalizability, particularly violating the second and third desiderata.
+Motivated by these limitations, we introduce an alternative approach inspired by the following key insights.
+First, many image restoration tasks can be intuitively described using natural language.
+In addition, large pre-trained text-to-image diffusion models like Stable Diffusion (Rombach et al., 2021) already capture rich distributions of both target images and common degradation artifacts described by these prompts.
+Based on these insights, we propose to approximate the joint prior’s score using large-scale pre-trained text-to-image diffusion models through classifier-free guidance (Ho and Salimans, 2021).
+In particular, we encode the desired target image characteristics using the positive prompts, and specify the unwanted degradation artifacts via a negative prompts.
+For example, when restoring a human face degraded by Gaussian blur, a suitable positive prompt can be “a clear headshot of a person”, and the corresponding negative prompt could be “blurry, low-quality image”.
+Formally, denoting the positive prompt as $\bm{c}$$_{+}$ and the negative prompt as $\bm{c}$$_{-}$, we can approximate $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ as
+$\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-})+\gamma(\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{+})-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-}))$ (6)
+where $\gamma>1$ is a weighting hyperparameter. When using latent diffusion models parameterized by $\theta$, this translates to the empirical noise prediction
+$\small\hat{\epsilon}_{\theta}({\bm{z}}_{t},t)=\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-})+\gamma(\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{+})-\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-}))$ (7)
+This straightforward approximation provides access to the otherwise intractable joint prior score, and enables versatile applications across diverse image and operator distributions.
+By leveraging the knowledge encoded in large pre-trained text-to-image models, our method bypasses the need for task-specific training, operator restriction or model re-selection.
+While this approach may appear to compromise the true "blindness", we emphasize that all blind inverse problem methods require some prior over the transformation space.
+Notably, textual descriptions of the target image and degradation artifacts are naturally advantageous, since users can directly infer the appropriate prompts from the measurements, which is a significantly simpler process than previous methods.
+These prompts provides an intuitive interface to encode diverse characteristics of both the desired outputs and the unwanted artifacts without training or handcrafted functions.
 Algorithm 1 Our algorithm LADiBI
-1: /* Initialize latent with encoded measurement ${\bm{y}}$ SDEdit */
-2:${\bm{z}}_{T_{s}}=\sqrt{\bar{\alpha}_{T_{s}}}{\mathbf{E}}(y)+\sqrt{1-\bar{\alpha}_{T_{s}}}\epsilon_{T_{s}}$ for $\epsilon_{T_{s}}\sim\mathcal{N}(0,I)$
+1: /* Initialize latent with encoded measurement $\bm{y}$} & SDEdit */
+[…]
 3:Initialize $\hat{\phi}$ with a fixed operator prior or Algorithm 2
@@
 10:   if $j\equiv 0\pmod{2}$ then
-11:              /* Perform MPGD with the estimated ${\mathcal{A}}_{\hat{\phi}}$/
+11:      /* Perform MPGD with the estimated $^{\mathcal{A}}_{\hat{\phi}}$ */
 12:     Update ${\bm{z}}_{0|t}$ with Eq. 8
 13:   end if
-14:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t},t)+\sigma_{t}\epsilon_{t}$ for $\epsilon_{t}\sim\mathcal{N}(0,I)$
-15:         ${\bm{z}}_{t}=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}{\bm{z}}_{t}+\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\epsilon^{\prime}_{t}$ for $\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I)$
-16:    end for
-17:     /* Periodically update $\hat{\phi}$ with ${\bm{z}}_{0|t}$ and Eq. 9 */
-18:    if $t\equiv 0\pmod{5}$ then
-19:         for $k=1,\dots,K$ do
-20:             $\hat{\phi}=\text{Adam}(\hat{\phi},L_{\phi}({\bm{z}}_{0|t},\hat{\phi}))$
-21:         end for
-22:    end if
-23:end for
-24:return ${\bm{x}}_{0}={\mathbf{D}}({\bm{z}}_{0})$
-3.1 Obtaining a better Prior
-Most existing diffusion-based methods [8, 27, 34, 23, 1, 39] avoid modeling the joint prior and instead further decompose the goal distribution as $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}})p({\mathcal{A}}_{\phi})$, under the assumption that the target data distribution $p({\bm{x}})$ independently from the operator distribution $p({\mathcal{A}}_{\phi})$.
-In this way, they can separately model the image and operator distribution by pre-trained diffusion models or fixed distributions and thus bypass the combinatorial complexity.
-However, is this truly a fair assumption?
-Consider the task of deblurring and super-resolution.
-A low-resolution blurry measurement may correspond to a high-resolution blurry image (e.g.
-an out of focus DSLR photo) if the operator is a downsampler, or to a low-resolution clear image if the operator is a blur kernel.
-Similarly, different target image distribution can imply different operators: for example, the estimated blur kernel should differ if the target is a realistic photo versus an Impressionist painting, even when both share the same measurement.
-These observations suggest that $p({\bm{x}})$ and $p({\mathcal{A}}_{\phi})$ should be correlated, and previous methods can make this independence assumption because they implicitly restrict the types of operations and target images they handle.
-In other words, while this independence assumption reduces modeling complexity, it also significantly limits the flexibility of the resulting algorithms.
-Consequently, most existing methods require additional training or parametrization of the operator prior and need to switch the image prior models when the target distribution changes.
-Therefore, to build a unified training-free algorithm, we should directly model $p({\bm{x}},{\mathcal{A}}_{\phi})$, or in the case of diffusion models, $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$.
-On the first glance, modeling $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ appears intractable.
-However, with large-scale pre-trained text-to-image diffusion models now widely available, we have a surprisingly simple solution:
-In practice, the target data distribution (e.g.
-“a high-resolution photo”) and the forward operation (e.g.
-“Gaussian blur”) can often be effectively described in natural language.
-We can specify the target distribution by using its description as a positive prompt.
-Specifying the operator distribution is less direct, as these pre-trained models are designed to represent image distributions rather than operators.
-However, by leveraging classifier-free guidance (CFG) [15], we can implicitly define the operator distribution through negative prompts.
-For example, if the operator is Gaussian blur, an effective negative prompt can be “blurry image, low quality”.
-Formally, denote positive prompt as ${\bm{c}}_{+}$ and negative prompt as ${\bm{c}}_{-}$, we can approximate $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ as
-$\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-})\\ +\gamma(\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{+})-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-}))$ (6)
-where $\gamma>1$ is a weighting hyperparameter.
-This is equivalent to approximating $p({\bm{x}},{\mathcal{A}}_{\phi})$ as a distribution that is proportional to $p({\bm{x}}|{\bm{c}}_{-})^{1-\gamma}p({\bm{x}}|{\bm{c}}_{+})^{\gamma}$.
-When using latent diffusion models parameterized by $\theta$, this corresponds to having the empirical noise prediction
-$\small\hat{\epsilon}_{\theta}({\bm{z}}_{t},t)=\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-})+\gamma(\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{+})-\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}}_{-}))$ (7)
-This way, our method not only models this seemingly extremely intractable joint distribution in a remarkably simple way, it also enable wide range of applications with diverse target image and operator distributions by leveraging the knowledge in large-scale pre-trained text-to-image diffusion models like Stable Diffusion[31].
-3.2 Obtaining a better Posterior
+14:   $\epsilon_{t},\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I)$
+15:   ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t},t)+\sigma_{t}\epsilon_{t}$
+16:   ${\bm{z}}_{t}=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}{\bm{z}}_{t-1}+\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\epsilon^{\prime}_{t}$
+17:  end for
+18:   /* Periodically update $\hat{\phi}$ with ${\bm{z}}_{0|t}$ and Eq. 9 */
+19:  if $t\equiv 0\pmod{5}$ then
+20:   for $k=1,\dots,K$ do
+21:     $)\hat{\phi}=\text{Adam}(\hat{\phi},L_{\phi}({\bm{z}}_{0|t},\hat{\phi}$))
+22:   end for
+23:  end if
+24:end for
+25:return ${\bm{x}}_{0}={\mathbf{D}}({\bm{z}}_{0})$
 Algorithm 2 General Operator Initialization
 1:for $j=1,\dots,M$ do
-2:     /* Initialize latent batch with encoded ${\bm{y}}$ SDEdit */
+2:   /* Initialize latent batch with encoded $\bm{y}$} & SDEdit */
 3:  $\epsilon^{(i)}_{T_{s}}\sim\mathcal{N}(0,I)$ for $i=1,2,\dots,N$
 4:  ${\bm{z}}_{T_{s}}^{(i)}=\sqrt{\bar{\alpha}_{T_{s}}}{\mathbf{E}}(y)+\sqrt{1-\bar{\alpha}_{T_{s}}}\epsilon^{(i)}_{T_{s}}$
-5:     /* Use MPGD to obtain latent estimations $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$/
+5:   /* Use MPGD to obtain latent estimations $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$ */
 6:  for $t=T_{j},\dots,1$ and all $i$ in parallel do
@@
 15:  for $k=1,\dots,K$ do
-16:         $\hat{\phi}=\text{Adam}(\hat{\phi},\frac{1}{N}\sum_{i=1}^{N}L_{\phi}({\bm{z}}_{0}^{(i)},\hat{\phi}))$
+16:   $)\hat{\phi}=\text{Adam}(\hat{\phi},\frac{1}{N}\sum_{i=1}^{N}L_{\phi}({\bm{z}}_{0}^{(i)},\hat{\phi}$))
 17:  end for
@@
 19:return $\hat{\phi}$
-In addition to a strong prior, our final output image should also satisfy the measurement constraint.Given the problem setup in Equation 4, the measurements are subject to additive Gaussian noise ${\bm{n}}$, hence $\log p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})=-\frac{1}{2\sigma^{2}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\bm{x}})\|^{2}_{2}$.
-When ${\mathcal{A}}_{\phi}$ is known, we can use the MPGD update rule in Equation 3 to perform posterior sampling.
-However, since ${\mathcal{A}}_{\phi}$ is not known, the true parameters $\phi$ is often approximated by another set of parameters $\hat{\phi}$.
+3.2 Sampling from the posterior
+In addition to a strong prior score, our output image should also satisfy the measurement constraint.
+Given the problem setup in Equation 4, the measurements are subject to additive Gaussian noise $\bm{n}$}, hence $\log p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})=-\frac{1}{2\sigma^{2}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\bm{x}})\|^{2}_{2}$.
+When ${\mathcal{A}}_{\phi}$ is known, we can use the MPGD update rule in Equation 3 for posterior sampling.
+However, since ${\mathcal{A}}_{\phi}$ is unknown, the true parameters $\phi$ is often approximated by another set of parameters $\hat{\phi}$.
 This approximation is usually addressed by one of the two strategies: an alternating optimization scheme that jointly approximates $\bm{x}$} and $\phi$, or obtaining a reliable $\hat{\phi}$ first then solving a non-blind inverse problem.
-The first approach is well-suited for training-free settings as it can be well integrated with diffusion-based methods, but it is often highly sensitive to $\hat{\phi}$ initialization and tuning.
-The second approach can perform well if a strong $\hat{\phi}$ is obtained, though it usually requires training and additional assumptions or restrictions.
-We propose to combine the two strategies above: we first obtain a reliable initial $\hat{\phi}$, and then perform an alternating optimization to iteratively refine both the operator parameter and data estimations throughout the diffusion process.
+The first approach is well-suited for training-free settings, but it is often highly sensitive to $\hat{\phi}$ initialization and tuning.
+The second approach can perform well if a strong $\hat{\phi}$ is obtained, though it usually requires training and additional restrictions.
+We propose a hybrid strategy: we first obtain a reliable initial $\hat{\phi}$, and then perform an alternating optimization to iteratively refine both the operator parameter and data estimations throughout the diffusion process.
 $\hat{\phi}$ Initialization
-In some cases, initializing the operator parameters is straightforward – for example, for a slight blur, an identity function can be a reasonable stating point.
-Alternatively, reliable operator priors, such as the pre-trained prior in BlindDPS [8] or the Gaussian prior in GibbsDDRM [27] for linear operators, can also provide effective initializations for simple kernels.
-However, when the operator requires complex functions like neural networks, obtaining good initializations becomes more challenging.
-To address this, we propose a new algorithm for general operator initialization by drawing further inspiration from alternating optimization algorithms.
-Note that since the goal here is only to initialize $\hat{\phi}$, the quality of intermediate ${\bm{x}}$ estimates is less critical as long as they provide a useful signal.
-Unlike other diffusion-based methods that alternate optimization targets at each diffusion step, we use SDEdit [26] and MPGD [14] with very few diffusion steps to quickly obtain a batch of ${\bm{x}}$ estimates.
+Sometimes initializing the operator parameters is straightforward – for example, for a slight blur, an identity function can be a reasonable starting point.
+Alternatively, reliable operator priors, such as the pre-trained prior in BlindDPS or the Gaussian prior in GibbsDDRM for linear operators, can also provide effective initializations for simple kernels.
+However, for complex operator estimator like neural networks, obtaining good initializations becomes more challenging.
+To address this, we propose a new algorithm for general operator initialization.
+Since the goal here is only to initialize $\hat{\phi}$, the quality of intermediate $\bm{x}$} estimates is less critical as long as they provide useful signals.
+Unlike other diffusion-based methods that alternate optimization targets at each diffusion step, we use SDEdit (Meng et al., 2022) and MPGD (He et al., 2024) with very few diffusion steps to quickly obtain a batch of $\bm{x}$} estimates.
 We then perform maximum likelihood estimation (MLE) on these estimates to update $\hat{\phi}$ using Adam optimizer.
 As detailed in Algorithm 2, repeating this process can leverage the diffusion prior to quickly converge to a reliable starting point for $\hat{\phi}$.
-Notice that we only assume that $\phi$ can be approximated by differentiable functions, allowing ${\mathcal{A}}_{\hat{\phi}}$ to be parameterized by general model families such as neural networks.
+Notice that we only assume that $\phi$ can be approximated by differentiable functions, allowing $^{\mathcal{A}}_{\hat{\phi}}$ to be parameterized by general model families such as neural networks.
 Moreover, as mentioned earlier, any well-performing operator priors can be seamlessly integrated into our framework.
 Iterative Refinement
-After initializing $\hat{\phi}$, we then perform another alternating optimization scheme to refine both the operator and the data.
-In particular, we alternate between updating ${\bm{z}}_{0|t}$ using MPGD guidance with ${\mathcal{A}}_{\hat{\phi}}$ fixed and updating the MLE estimation of $\hat{\phi}$ using Adam with ${\bm{z}}_{0|t}$ fixed throughout the diffusion process.
+After $\hat{\phi}$ initialization, we perform another alternating optimization to refine both the operator and the image.
+Throughout the diffusion process, we alternate between updating ${\bm{z}}_{0|t}$ using MPGD with $^{\mathcal{A}}_{\hat{\phi}}$ fixed and updating the MLE estimation of $\hat{\phi}$ using Adam with ${\bm{z}}_{0|t}$ fixed.
 Since unlike GibbsDDRM, which uses Langevin dynamics to update the operator, we solve for a local optimum.
 Therefore, it’s reasonable not to update the operator too frequently.
-Empirically, we find that periodic updates to the operator in combination of the time-traveling technique [25, 42] yield the best results
-In addition, since MPGD supports any differentiable loss function, we can incorporate techniques like regularization to further improve the visual quality. In practice, we use
+Empirically, we find that periodic updates to the operator combined with time-traveling (Lugmayr et al., 2022; Yu et al., 2023) yield the best results.
+In addition, since MPGD supports any differentiable loss function, we can incorporate regularization to further improve the visual quality.
+In practice, we use
 $\small{\bm{z}}_{0|t}={\bm{z}}_{0|t}-c_{t}\nabla_{{\bm{z}}_{0|t}}(\|{\bm{y}}-{\mathcal{A}}_{\hat{\phi}}({\mathbf{D}}({\bm{z}}_{0|t}))\|^{2}_{2}+\lambda_{\bm{z}}\text{LPIPS}({\mathbf{D}}({\bm{z}}_{0|t}),{\bm{y}}))$ (8)
@@
 $\lambda_{\bm{z}}$ and $\lambda_{\phi}$ are adjustable hyperparameters.
-By combining large-scale language-conditioned diffusion priors, effective operator initializations, and the iterative refinement process described above, we introduce LADiBI (Algorithm 1), a new training-free algorithm for blind inverse problem solving that supports diverse target image distributions and flexible operators.
+Combining large-scale language-conditioned diffusion priors, effective operator initializations and the iterative refinement process described above, we introduce LADiBI, a new training-free algorithm for blind inverse problem solving that supports diverse target image distributions and flexible operators.
 4 Experiments
@@
 We empirically verify the performance of our method with two linear deblurring tasks, Gaussian deblurring and motion deblurring, and a non-linear restoration task, JPEG decompression.
-Here we provide the details of our experiments.
-Datasets
-We conduct quantitative comparisons on the datasets FFHQ 256 $\times$ 256 [18] and AFHQ-dog 256 $\times$ 256 [6].
-We use a size of 1000 images for FFHQ and 500 for AFHQ, following the setting in Murata et al.
-[27].
+Setup
+We conduct quantitative comparisons on FFHQ 256 $\times$ 256 (Karras et al., 2019) and AFHQ-dog 256 $\times$ 256 (Choi et al., 2020).
+Following Murata et al.
+(2023), we use 1000 images for FFHQ and 500 images for AFHQ.
+We use three quantitative metrics: LPIPS (Zhang et al., 2018), PSNR and KID (Bińkowski et al., 2018).
+PSNR directly measures the pixel-by-pixel accuracy of the reconstructed image compared to the original, while LPIPS measures how similar two images are in a way that reflects human perception.
+Finally, KID evaluates the image fidelity on a distribution level.
 Baselines
 We compare our method against other state-of-the-art approaches as baselines.
-Specifically, we choose Pan-l0 [28] and Pan-DCP [29] as the optimization-based method, SelfDeblur [30] as the self-supervised approach, PRNet [44] and DeblurGANv2 [22] as the supervised baselines, and BlindDPS [8] and GibbsDDRM [27] as the diffusion-based methods.
-All baselines are experiments using their open-sourced code and pre-trained models, whose details are provided in the appendix.
+Specifically, we choose Pan-l0 (Pan et al., 2017) and Pan-DCP (Pan et al., 2018) as the optimization-based method, SelfDeblur (Ren et al., 2020) as the self-supervised approach, PRNet (Zamir et al., 2021) and DeblurGANv2 (Kupyn et al., 2019) as the supervised baselines, and BlindDPS (Chung et al., 2023a), BIRD (Chihaoui et al., 2024) and GibbsDDRM (Murata et al., 2023) as the diffusion-based methods.
+All baselines are experiments using their open-sourced code and pre-trained models.
 Implementation Details
-We conduct all experiments an NVIDIA A6000 GPU.
-We use Stable Diffusion 1.4 [31], DDIM $200$-step noise scheduling, and $T_{s}=150,M=4$ for all tasks.
-We set positive prompts as “a clear headshot of a person” and “a clear headshot of an animal” for FFHQ and AFHQ experiments respectively.
-We adjust negative prompts according to each individual task, and the details will be provided in the following sections.
-Evaluation Metrics
-We evaluate our method and the baselines with three quantitative metrics: LPIPS [46], PSNR and KID [4].
-Each metric has its own significance and reveals a different set of characteristics about performance.
-In particular, PSNR directly measures the pixel-by-pixel accuracy of the reconstructed image compared to the original, while LPIPS measures how similar two images are in a way that reflects human perception.
-Finally, KID evaluates the image fidelity on a distribution level.
-4.2 Blind Linear Deblurring
-We first test our method against the baselines on linear deblurring, which is the design space of most baselines. We evaluate all methods on two types of blur kernels: Gaussian blur and motion blur.
-Setup
-To generate measurements, we apply randomly generated motion blur kernels with intensity 0.5 and Gaussian blur kernels with standard deviation 3.0 respectively.
-The final measurements are derived by applying a pixel-wise Gaussian noise with $\sigma=0.02$.
-We use a $61\times 61$ convolutional matrix as $\hat{\phi}$ and we initialize it as a Gaussian kernel of intensity 6.0 following Murata et al.
-[27].
-We adjust the negative prompts to the potential negative image artifacts induced by the degradation operation, such as “lowres, blurry, DSLR effect, poorly drawn face, deformed, bad proportions, …” for Gaussian blur, “shaken image, morbid, mutilated, text in image, disfigured, …” for motion blur.
-Full prompts are included in the appendix.
-Results
+We conduct all experiments on an NVIDIA A6000 GPU.
+We use Stable Diffusion 1.4 (Rombach et al., 2021), DDIM $200$-step, and $T_{s}=150,M=4$ for all tasks.
+We set positive prompts as “a clear headshot of a person" and “a clear headshot of an animal" for FFHQ and AFHQ experiments respectively.
+We adjust negative prompts according to each individual task and the details are be provided in the Appendix A.
 Table 2: Quantitative results on blind linear deblurring tasks.
@@
 […]
+DPS w/ GT kernel 0.164 22.82 0.0046 0.138 24.48 0.0052 0.367 21.28 0.1120 0.330 23.52 0.0813
 SelfDeblur 0.732 9.05 0.1088 0.733 8.87 0.0890 0.742 9.04 0.0650 0.736 8.84 0.0352
@@
 BlindDPS 0.246 20.93 0.0153 0.216 25.96 0.0205 0.393 20.14 0.0913 0.330 24.79 0.0268
+BIRD 0.294 19.23 0.0491 0.212 21.95 0.0414 0.438 18.92 0.0286 0.320 21.87 0.0089
 GibbsDDRM 0.293 20.52 0.0746 0.216 27.03 0.0430 0.303 19.44 0.0265 0.257 24.01 0.0040
@@
 From top to bottom we showcase examples from motion deblur on FFHQ, Gaussian deblur on FFHQ, motion deblur on AFHQ, and Gaussian deblur on AFHQ respectively.
-Table 2 and Figure 2 show results on blind linear deblurring.
-Our method matches the performance the state-of-the-art method GibbsDDRM, which is explicitly designed to solving linear problems using SVDs.
+4.2 Blind Linear Deblurring
+We first conduct experiments on linear deblurring, which is the design space of most baselines.
+We evaluate all methods on two blurring kernels: Gaussian blur and motion blur.
+We apply random motion blur kernels with intensity $0.5$ and Gaussian blur kernels with standard deviation $3$.
+Measurements are derived by applying a pixel-wise Gaussian noise with $\sigma=0.02$.
+We use a $61\times 61$ convolutional matrix as $\hat{\phi}$ and initialize it as a Gaussian kernel of intensity $6.0$ following Murata et al.
+(2023)
+As shown in Table 2 and Figure 3, our method matches the performance of the state-of-the-art method GibbsDDRM, which is explicitly designed to solve linear problems using SVDs.
 In fact, GibbsDDRM’s highly specialized design makes it so sensitive that even small discrepancies (e.g.
@@
 Moreover, although BlindDPS and GibbsDDRM use diffusion models that are trained on the exact data distributions tested, our method can still match or outperform them with the large-scale pre-trained model.
-In addition, even though supervised methods can obtain higher PSNR, our method produces the fewest artifacts, which is also validated by the LPIPS and KID scores.
-Overall, our method offers a consistently competitive performance on linear tasks, even though our method is designed for more general applications.
+While supervised methods obtain higher PSNR, our method produces the fewest artifacts, validated by LPIPS and KID scores.
+Overall, our method offers a competitive performance on linear tasks, even though it is designed for more general applications.
 4.3 Blind JPEG Decompression
-We further compare our method with baselines on JPEG decompression, a particularly challenging task due to the non-linear, non-differentiable nature of JPEG compression.
+We further compare all methods on JPEG decompression, a particularly challenging task due to its non-linear, non-differentiable nature.
 Unlike traditional settings, our experiments provide no task-specific knowledge, such as the compression algorithm, quantization table, or factors – algorithms rely solely on the measurement images.
-Setup
 We generate measurements using JPEG compression with a quantization factor of 2.
-We use a neural network with a 3-layer U-net [32] to parametrize the operator.
+We use a neural network with a 3-layer U-net (Ronneberger et al., 2015) to parametrize the operator.
 The operator is initialized using Algorithm 2 with $M=8$ and $N=4$.
-We use phrases including “pixelated, low quality, jpeg artifacts, deformed, bad proportions, …” as negative prompt, whose full version is provided in the appendix.
-Results
-Table 3: Quantitative results on blind JPEG decompression task.
+Table 3: Quantitative results (the mean and standard deviation) on blind JPEG decompression task.
 Method FFHQ AFHQ
 LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
-SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
-MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
-DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
-BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
-GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
-LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
+Pan-I0 0.787 $\pm$ 0.084 12.72 $\pm$ 2.53 0.2189 0.825 $\pm$ 0.083 13.15 $\pm$ 2.68 0.2168
+Pan-DCP 0.710 $\pm$ 0.067 14.91 $\pm$ 2.63 0.2318 0.673 $\pm$ 0.055 18.27 $\pm$ 1.60 0.2472
+SelfDeblur 0.676 $\pm$ 0.054 8.84 $\pm$ 1.93 0.1959 0.703 $\pm$ 0.052 8.89 $\pm$ 1.72 0.1173
+MPRNet 0.785 $\pm$ 0.048 5.99 $\pm$ 1.81 0.8008 0.769 $\pm$ 0.048 5.93 $\pm$ 1.78 0.6809
+DeblurGAN 0.473 $\pm$ 0.058 21.48 $\pm$ 1.85 0.2207 0.502 $\pm$ 0.061 21.76 $\pm$ 1.84 0.1527
+BlindDPS 0.431 $\pm$ 0.079 21.55 $\pm$ 1.96 0.1791 0.397 $\pm$ 0.069 20.87 $\pm$ 1.89 0.2108
+BIRD 0.406 $\pm$ 0.047 20.68 $\pm$ 1.09 0.0525 0.425 $\pm$ 0.068 21.08 $\pm$ 1.50 0.0673
+GibbsDDRM 0.841 $\pm$ 0.057 13.91 $\pm$ 1.26 0.2915 0.775 $\pm$ 0.059 14.59 $\pm$ 1.48 0.2915
+Ours 0.268 $\pm$ 0.070 21.40 $\pm$ 1.18 0.0172 0.315 $\pm$ 0.075 21.12 $\pm$ 1.27 0.0216
 Figure 4: Qualitative results on the blind JPEG decompression task.
@@
 Unlike the baselines, which struggle with this task due to their limited posterior formulations, our flexible framework adapts to approximate this challenging operator and produces high quality images.
-4.4 Ablation Studies
-Table 4: Ablation study on our proposed algorithm.
-Removed Component LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
-Text prompts 0.500 18.88 0.0483
-MPGD guidance 0.470 17.80 0.0434
-$\hat{\phi}$ Initialization 0.355 19.74 0.0250
-LADiBI 0.315 21.12 0.0216
+4.4 Ablation Study
+Table 4: Ablation study on JPEG decompression.
+Ablation LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
+W/o text prompts 0.508 19.30 0.0319
+W/o negative prompts 0.440 19.44 0.0242
+Using generic $\bm{c}$$_{-}$ 0.425 19.48 0.0255
+Using task-irrelevant $\bm{c}$$_{-}$ 0.406 19.40 0.0237
+W/o MPGD guidance 0.403 19.44 0.0274
+W/o regularization 0.307 21.35 0.0172
+W/o $\hat{\phi}$ Initialization 0.289 20.17 0.0170
+LADiBI 0.262 21.20 0.0132
+Figure 5: Qualitative results on blind deblurring Monet paintings.
 To showcase the effectiveness of each part of our algorithm, we conduct ablation study on AFHQ dataset with JPEG decompression task.
-In particular, we test the importance of suitable ad-hoc prompts, the use of MPGD guidance, and the operator initialization for neural network ${\mathcal{A}}_{\hat{\phi}}$.
-in-between diffusion steps.
+In particular, we test the importance of suitable ad-hoc prompts, the use of MPGD guidance, the MPGD regularization term, and the operator initialization for neural network $^{\mathcal{A}}_{\hat{\phi}}$.
 We keep SDEdit in all settings to encode the measurement information for fair comparisons.
@@
 4.5 Style Preserving Deblurring
-Figure 4: Qualitative results on blind deblurring Monet paintings.
-In order to demonstrate our method’s ability to solve blind inverse problems from a wider range of image distributions, we present indicative examples of Gaussian and motion deblurring on Monet paintings in Figure 4.
-We use the same negative prompts as in the previous experiments and “a portrait of a person as a Monet style painting” as positive prompts.
+In order to show our method’s applicability on a wider range of image and operator distributions, we present indicative examples of Gaussian and motion deblurring on Monet paintings in Figure 5.
+Additional operators and image distributions demonstrations are available in Appendix B.
+We use the same negative prompts as in the previous experiments and “a portrait of a person as a Monet style painting" as positive prompts.
 We notice that, although the images portray human faces, the baseline method using models trained on realistic human face images cannot accurately solve the problem, while our method effectively generates images consistent with both the measurement image and the painting style present in the ground truth image.
 5 Conclusion
-In this work, we propose LADiBI, a new training-free algorithm to solving blind inverse problems in image restoration using a large-scale pre-trained text-to-image diffusion model.
-Our method leverages text prompts as well as posterior guidance on intermediate diffusion steps to guide the generation process towards clean reconstructions based on degraded measurements, in cases where the degradation operation is unknown.
-Experimental results demonstrate that LADiBI achieves competitive performance on tasks with diverse operator and is also capable of covering a wider range of image distributions compared to baselines.
+In this work, we propose LADiBI, a new training-free algorithm to solving blind inverse problems in image restoration using large-scale pre-trained text-to-image diffusion models.
+With unknown degradation operators, our method leverages text prompts as well as posterior guidance on intermediate diffusion steps to restore desired images based on the measurements.
+Experiments demonstrate that LADiBI’s effectiveness on diverse operator and image distributions.
 Acknowledgment
@@
 Appendix A Additional Implementation Details
-We offer more detailed information regarding the implementation setup for Algorithms 1 and 2.
+In this section, we offer more detailed information regarding the implementation setup for Algorithms 1 and 2.
 All experiments are conducted using an NVIDIA A6000 GPU.
@@
 In parallel with the reverse diffusion process, we update $\hat{\phi}$ every 5 timesteps, each time conducting $K=150$ gradient steps.
-We adust the $c_{t}$ and the $\lambda_{\phi}$ parameters of Equations 8 and 9 to 30 and 2 respectively.
+We adjust the $c_{t}$ and the $\lambda_{\phi}$ parameters of Equations 8 and 9 to 30 and 2 respectively.
 In terms of the operator initialization process described by Algorithm 2, we make use of a batch of $N=4$ samples and run $M=8$ iterations, each comprising $T_{j}=60$ timesteps.
-Figure 5: A schematic overview of Algorithm 1.
-Figure 6: A schematic overview of Algorithm 2.
-Additionally, schematic overviews of Algorithms 1 and 2 are presented in Figures 5 and 6.
-Furthermore, there are some parameters in our approach for which employing a task-aware setup strategy is essential for state-of-the-art performance.
+We also take into consideration that the targets and measurements reside in the $256\times 256$ pixel space, whereas Stable Diffusion v1 operates on images of pixel size $512\times 512$.
+To address this disparity, we initially upsample the measurement using bilinear interpolation in order to transform it to a $512\times 512$ image, and then downsample the resulting $x_{0}$ to map the final estimate back to the original image space.
+The schematic overviews of Algorithms 1 and 2 are presented in Figures 2 and 6.
+In addition, there are some parameters in our approach for which employing a task-aware setup strategy is essential for state-of-the-art performance.
 These are the prompt sentences as well as the architecture of $\hat{{\phi}}$.
@@
 Motion Deblurring
-  • Positive prompts: “a clear headshot of a person/animal”
-  • Negative prompts: “shaken image, motion blur, pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
+  • Positive prompts: “a clear headshot of a person/animal"
+  • Negative prompts: “shaken image, motion blur, pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
   • $\hat{\phi}$ architecture: A single $61\times 61$ convolutional block with 3 input and 3 output channels.
 Gaussian Deblurring
-  • Positive prompts: “a clear headshot of a person/animal”
-  • Negative prompts: “blurry, gaussian blur, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
+  • Positive prompts: “a clear headshot of a person/animal"
+  • Negative prompts: “blurry, gaussian blur, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
   • $\hat{\phi}$ architecture: A single $61\times 61$ convolutional block with 3 input and 3 output channels.
 JPEG Decompression
-  • Positive prompts: “a clear headshot of a person/animal”
-  • Negative prompts: “pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, jpeg artifacts, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
-  • $\hat{\phi}$ architecture: A neural network with a typical 3-layer U-net [32] architecture.
+  • Positive prompts: “a clear headshot of a person/animal"
+  • Negative prompts: “pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, jpeg artifacts, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
+  • $\hat{\phi}$ architecture: A neural network with a typical 3-layer U-net [Ronneberger et al., 2015] architecture.
 Each layer consists of 2 convolutional blocks of size $3\times 3$ with ReLU activations and number of input and output channels ranging from 32 to 128.
+We conduct quantitative experiments on FFHQ $256\times 256$ [Karras et al., 2019] and AFHQ $256\times 256$ [Choi et al., 2020].
+Images in FFHQ are publicized under Creative Commons BY 2.0, Creative Commons BY-NC 2.0, Public Domain Mark 1.0, Public Domain CC0 1.0, or U.S.
+Government Works license and AFHQ is publicized under Attribution-NonCommercial 4.0 International license.
+Figure 6: A schematic overview of general operator initialization (Algorithm 2).
 Appendix B Additional Results
-Table 5: Additional quantitative results on blind JPEG decompression task.
-Method FFHQ AFHQ
+In this section, we present additional experimental results on all benchmarks as well as additional tasks and data distributions.
+Table 5 include the standard deviation of the linear benchmark results.
+Figures 7, 8 and 9 offer more qualitative comparisons against baseline methods on motion deblurring, gaussian deblurring and JPEG decompression respectively.
+In Figure 10, we include comparisons against selected baseline methods on the 3 benchmark tasks using a set of Monet and Van Gogh portrait paintings as ground truth images.
+We also show additional demonstrations of applying LADiBI to solve colorization and non-linear deblurring problems.
+Figures 12 and 12 substantiate our claim that our algorithmic scheme is flexible enough to adapt to various inverse tasks.
+Additionally, thanks to the broad range of image distributions encapsulated by the pretrained model, as well as the minimal assumptions imposed by the proposed methodology, our approach is able to accurately solve inverse problems for a broad range of images.
+Figures 14 and 14 demonstrate our method’s capability to perform Gaussian deblurring, motion deblurring as well as JPEG decompression for images depicting cars and landscapes.
+Finally, to assess the real-world performance of our method, we test it on samples from the RealBlur dataset [Rim et al., 2020], which contains naturally captured blurry images using GoPro in real life settings.
+This dataset presents realistic blur patterns that are significantly more complex and diverse compared to the synthetically generated benchmarks.
+Figure 15 presents qualitative demos, which highlight our method’s ability to recover sharp structures and textures under various real-world blur conditions.
+These results demonstrate the applicability of our approach to real inverse problem solving tasks and support its generalization capability.
+Table 5: Quantitative results on motion and Gaussian deblurring on FFHQ dataset. Mean $\pm$ standard deviation shown.
+Method Motion Gaussian
 LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
-Pan-l0 0.787 12.72 0.2190 0.825 13.15 0.2168
-Pan-DCP 0.710 14.91 0.2318 0.673 18.27 0.2472
-SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
-MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
-DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
-BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
-GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
-LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
+DPS w/ GT kernel 0.164 22.82 0.0046 0.138 24.48 0.0052
+SelfDeblur $0.732\pm 0.147$ $9.05\pm 2.46$ 0.1088 $0.733\pm 0.094$ $8.87\pm 2.21$ 0.0890
+MPRNet $0.292\pm 0.101$ $\underline{22.42}\pm 3.21$ 0.0467 $0.334\pm 0.068$ $23.23\pm 2.19$ 0.0511
+DeblurGANv2 $0.309\pm 0.111$ $\mathbf{22.55}\pm 3.44$ 0.0411 $0.325\pm 0.145$ $26.61\pm 3.28$ 0.0227
+Pan-l0 $0.389\pm 0.090$ $14.10\pm 2.51$ 0.1961 $0.265\pm 0.082$ $20.68\pm 3.81$ 0.1012
+Pan-DCP $0.325\pm 0.105$ $17.64\pm 3.63$ 0.1323 $0.235\pm 0.066$ $24.93\pm 3.60$ 0.0490
+BlindDPS $0.246\pm 0.077$ $20.93\pm 2.09$ 0.0153 $0.216\pm 0.076$ $25.96\pm 2.45$ 0.0205
+BIRD $0.294\pm 0.076$ $19.23\pm 1.88$ 0.0491 $0.212\pm 0.055$ $21.95\pm 1.62$ 0.0414
+GibbsDDRM $0.293\pm 0.099$ $20.52\pm 2.81$ 0.0746 $0.216\pm 0.046$ $\underline{27.03}\pm 1.87$ 0.0430
+LADiBI (Ours) $\underline{0.230}\pm 0.076$ $20.96\pm 2.34$ $\mathbf{0.0084}$ $\underline{0.197}\pm 0.071$ $21.08\pm 2.71$ $\mathbf{0.0068}$
+GibbsDDRM* $\textbf{0.199}\pm 0.110$ $22.36\pm 3.79$ 0.0309 $\textbf{0.155}\pm 0.049$ $\textbf{27.65}\pm 2.66$ 0.0252
+Table 6: Quantitative results on motion and Gaussian deblurring on AFHQ. Mean $\pm$ standard deviation shown.
+Method Motion Gaussian
+LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
+DPS w/ GT kernel 0.367 21.28 0.1120 0.330 23.52 0.0813
+SelfDeblur $0.742\pm 0.158$ $9.04\pm 1.84$ 0.0650 $0.736\pm 0.112$ $8.84\pm 1.45$ 0.0352
+MPRNet $0.324\pm 0.095$ $\underline{22.09}\pm 3.01$ 0.0382 $0.379\pm 0.081$ $21.97\pm 2.55$ 0.0461
+DeblurGANv2 $0.323\pm 0.105$ $\mathbf{22.74}\pm 2.89$ 0.0350 $0.340\pm 0.084$ $\mathbf{27.12}\pm 2.94$ 0.0073
+Pan-l0 $0.414\pm 0.133$ $14.16\pm 3.97$ 0.1590 $0.276\pm 0.079$ $21.04\pm 3.39$ 0.0320
+Pan-DCP $0.371\pm 0.147$ $17.63\pm 5.94$ 0.1377 $0.297\pm 0.086$ $\underline{25.11}\pm 3.68$ 0.0263
+BlindDPS $0.393\pm 0.061$ $20.14\pm 1.67$ 0.0913 $0.330\pm 0.057$ $24.79\pm 1.76$ 0.0268
+BIRD $0.438\pm 0.110$ $18.92\pm 2.04$ 0.0286 $0.320\pm 0.079$ $21.87\pm 1.95$ 0.0089
+GibbsDDRM $0.303\pm 0.114$ $19.44\pm 3.58$ 0.0265 $0.257\pm 0.121$ $24.01\pm 4.77$ $\mathbf{0.0040}$
+LADiBI (Ours) $\mathbf{0.262}\pm 0.096$ $21.20\pm 2.72$ $\mathbf{0.0132}$ $\mathbf{0.204}\pm 0.066$ $24.33\pm 1.73$ 0.0065
+GibbsDDRM* $\underline{0.278}\pm 0.099$ $19.00\pm 3.09$ 0.0180 $\underline{0.224}\pm 0.091$ $21.62\pm 3.53$ 0.0040
 Figure 7: Additional qualitative results on motion deblurring task.
@@
 Figure 10: Additional qualitative results on painting restoration.
-We present more detailed experimental results on all benchmark tasks.
-In specific, Table 5 shows quantitative results on the JPEG decompression task, including the optimization-based Pan-l0 and Pan-DCP methods.
-Moreover, Figures 7, 8 and 9 offer more qualitative comparisons against baseline methods on motion deblurring, gaussian deblurring and JPEG decompression respectively.
-Finally, in Fig.
-10 we include comparisons against selected baseline methods on the 3 benchmark tasks using a set of Monet and Van Gogh portrait paintings as ground truth images.
+Figure 11: Qualitative results on colorization task.
+Figure 12: Qualitative results on non-linear deblurring task.
+Figure 13: Demo inverse problem solving on car images.
+Figure 14: Demo inverse problem solving on landscape images.
+Figure 15: Qualitative samples on the RealBlur dataset.
 Appendix C Additional Ablation Study
@@
 Figure 17: Ablation study on number of time-traveling steps.
+Table 7: Quantitative comparison between linear and deep operator architecture
+AFHQ Motion Blur
+Method LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
+Panl0 0.414 14.16 0.1590
+PanDCP 0.371 17.63 0.1377
+SelfDeblur 0.742 9.04 0.0650
+MPRNet 0.324 22.09 0.0382
+DeblurGANv2 0.323 22.74 0.0350
+BlindDPS 0.393 20.14 0.0913
+GibbsDDRM 0.303 19.44 0.0265
+GibbsDDRM* 0.278 19.00 0.0180
+LADiBI (Kernel Operator) 0.262 21.20 0.0132
+LADiBI (U-Net Operator) 0.343 18.93 0.0146
 Figure 18: Comparison between linear and deep operator architecture on linear tasks.
-We carefully adjust each parameter of our methodology to the value that offers the best overall results by conducting comparative experiments and in this section we offer some indicative examples of such tests.
-All hyper-parameter tuning experiments have been performed on the motion deblurring task using the AFHQ dataset.
-Figure 11 presents both the LPIPS and KID score for the value of the timestep at which we encode the measurement using the forward diffusion process.
-Both metrics form a U-shape curve, which results from the following fact about encoding the measurement image; if we make the encoding timestep too large, most of the information about the measurement has been replaced by noise which does not allow the sampling process to leverage useful features of the degraded image.
-On the other hand, if the timestep is too small, the sampling process is not equipped with enough scheduled variance to reach the target prior distribution.
+In this section, we provide additional ablation study on various hyperparameters and components of our LADiBI algorithm.
+All ablation studies are conducted on the motion deblurring task using the AFHQ dataset.
+Figure 17 presents both the LPIPS and KID score for the value of the timestep $T_{s}$ at which we encode the measurement using SDEdit [Meng et al., 2022].
+Our experiments align with the observation in Meng et al.
+[2022]: If $T_{s}$ is too large most of the information about the measurement has been replaced by noise which does not allow the sampling process to leverage useful features of the degraded image.
+On the other hand, if $T_{s}$ is too small, the sampling process is not equipped with enough scheduled steps to reach the target distribution.
 In addition, Figure 17 presents the effectiveness of taking advantage of the time-traveling strategy.
 More time-traveling boosts the overall performance up to a specific value, after which we begin to notice a trade-off between perceptual clarity of the image and fidelity to the target distribution.
-With respect to ablation studies, we are interested in evaluating the performance of our algorithm when using a task-agnostic implementation for ${\mathcal{A}}_{\hat{\phi}}$.
-Hence, we tested the neural network architecture employed for non-linear tasks at linear inverse problems and we showcase qualitative results in Figure 13.
-We observe that, although performing worse than LADiBI employed with a linear operator architecture, the task-agnostic method is capable of producing estimates of decent quality while also preserving a design that allows applicability to any image restoration task.
-Appendix D Limitations and Future Works
+Finally, we evaluate the performance of our algorithm when using a unconstrained configuration for $^{\mathcal{A}}_{\hat{\phi}}$ in constrained tasks.
+In particular, we test the neural network architecture for the operator in linear inverse problems and we showcase quantitative results for the motion deblurring task in Table 7, and qualitative samples in Figure 18.
+We observe that, although performing worse than LADiBI employed with the aligned linear operator architecture, the neural operator is still capable of producing estimates of decent quality while also preserving a design that allows applicability to highly unconstrained image restoration task.
+Appendix D Prior Coverage of Large Pre-trained Text-to-image Diffusion
+In this section we provide an empirical justification for using Stable Diffusion v1.4 as our base model.
+To demonstrate this property, Figure 19 provides qualitative samples of images with Gaussian blur, motion blur and after JPEG compression.
+These examples show that Stable Diffusion 1.4 already capture rich distributions of both target images and common degradation artifacts.
+(a) Generated images with Gaussian blur effect.
+(b) Generated images with motion blur effect.
+(c) Generated images with JPEG compression effect.
+Figure 19: Drawn samples from the posterior distribution using our baseline model.
+Appendix E Limitations and Future Works
 In this section, we discuss the limitation of LADiBI and potential future works to address these problems.
-Similar to many training-free diffusion posterior sampling algorithms [8, 27, 9, 14], our method is also sensitive to hyper-parameter tuning.
+Similar to many training-free diffusion posterior sampling algorithms [Chung et al., 2023a, Murata et al., 2023, Chung et al., 2023b, He et al., 2024], our method is also sensitive to hyper-parameter tuning.
 We have provided details about our hyperparameter choices in the previous sections, and we will release our code in a public repository upon publication of this paper.
+In addition, while generally simple, we do require the users to infer appropriate prompts from the measurement.
+An interesting future work direction can include automated prompt tuning similar to the method proposed in Chung et al.
+[2023c].
 Another notable drawback of our method is that LADiBI requires significantly longer inference time in comparison to the best performing baseline (i.e.
-GibbsDDRM [27]) in order to obtain high quality restorations.
+GibbsDDRM [Murata et al., 2023]) in order to obtain high quality restorations.
 With the general operator initialization, our algorithm can take around 5 minutes to complete on a single NVIDIA A6000 GPU while GibbsDDRM only takes around 30 seconds.
@@
 Exploring neural network architectures that can easily generalize across different operator functional classes while achieving state-of-the-art results efficiently remains an exciting direction for future work.
-Appendix E Broader Impact Statement
+Appendix F Impact Statement
 Lastly, since our algorithm leverages large pre-trained image generative models, we would like to address the ethical concerns, societal impact as well as the potential harm that can be caused by our method when being used inappropriately.